### v2

# Optimal payoff under Bregman-Wasserstein divergence constraintsthanks: Second version: October 28, 2025. First version November 21, 2024

Abstract
We study optimal payoff choice for an expected utility maximizer under the constraint that their payoff is not allowed to deviate “too much” from a given benchmark. We solve this problem when the deviation is assessed via a Bregman-Wasserstein (BW) divergence, generated by a convex function $\phi$. Unlike the Wasserstein distance (i.e., when $\phi(x)=x^{2}$) the inherent asymmetry of the BW divergence makes it possible to penalize positive deviations different than negative ones. As a main contribution, we provide the optimal payoff in this setting. Numerical examples illustrate that the choice of $\phi$ allow to better align the payoff choice with the objectives of investors.
keywords:
Portfolio choice, Preferences, Expected Utility, Hoeffding–Fréchet bounds, Bregman Divergence, Wasserstein distance.
1 Introduction
A standard assumption in the literature on optimal payoff choice is that investors’ preferences can be described by the expected utility theory (EUT) of von Neumann and Morgenstern, (1947). This approach has however faced significant criticisms and researchers have explored optimal payoff choices under alternative decision theories. Notable are the dual theory of Yaari, (1987), the rank-dependent expected utility (RDEUT) approach of Quiggin, (1993), and several behavioural approaches, including the SP/A theory of Lopes, (1987) and Shefrin and Statman, (2000), and the cumulative prospect theory (CPT) of Tversky and Kahneman, (1992). Despite extensive research within these alternative frameworks [^1] Optimal payoffs have been derived for an RDEUT investor in Carlier and Dana, (2011), Xia and Zhou, (2016), He and Zhou, (2016) Xu, (2016), Jaimungal et al., (2022); for a CPT investor in Jin and Zhou, (2008), Zhang et al., (2011), and Rüschendorf and Vanduffel, (2020); and for a Yaari investor in He and Zhou, (2011), He and Jiang, (2021) and Boudt et al., (2022)., none prove superior to EUT for the problem of optimal portfolio choice. For instance, under Yaari’s dual theory, optimal payoffs often materialize as binary (digital) or trinary options, which are infrequently practical since investors rarely opt for such payoffs. In addition, Bernard et al., 2015a demonstrate that the optimal payoff of an investor with law-invariant increasing preferences can always be rationalized via a concave utility function, and their optimal payoff obtained by maximizing the corresponding expected utility.
In real-world investment contexts, investors commonly encounter constraints on fund allocation. For instance, shareholders of an investment company often seek a balance between risk and return, whereas a supervisory authority of a pension fund prioritizes the fund’s ability to meet participant obligations. Additionally, asset managers frequently establish a benchmark when engaging with clients, striving to outperform it without significant deviation. Moreover, within Markowitz’s seminal mean-variance framework, additional constraints are typically imposed on the composition of buy-and-hold portfolios to ensure “reasonable” allocations. As a consequence, various studies have delved into optimal payoff selection under diverse constraints. However, the majority of these works predominantly tackle distributional constraints of the payoff itself, rather than exploring the behaviour of payoffs in relation to benchmarks. For instance, Basak and Shapiro, (2001) derive the optimal payoff for an EUT maximizer under the constraint of maintaining the Value-at-Risk (VaR) within acceptable limits. Other studies in this realm include Wei, (2021), Cuoco et al., (2008), and Chen et al., (2024).
In this paper, we examine the optimal choices for an EUT investor who is concerned about the deviation of their payoff from a specified benchmark. Related problems has been investigated in Pesenti and Jaimungal, (2023) and Jaimungal et al., (2022). The former considers Yaari investors and used (among other constraints) the 2-Wasserstein distance to measure divergence while the latter focuses on RDEUT and the $p$-Wasserstein distance. The $p$-Wasserstein distance arises as the solution to the Monge-Kantorovich optimal transport (OT) problem with the symmetric cost function $c(z_{1},z_{2})=|z_{1}-z_{2}|^{p}$. Thus positive deviations (gains) from a benchmark are penalised to the same extent as negative ones (losses) [^2] Other research that has addressed optimal payoff choices under symmetric distance constraints include works by Bernard et al., 2015b , Rüschendorf and Vanduffel, (2020), and He and Jiang, (2021).. In optimal payoff choice, however, asymmetry may be desired. Specifically, since the fundamental work of Tversky and Kahneman, (1992), it is well accepted that individuals tend to feel the pain of losses more acutely than the pleasure from equivalent gains. We contribute to the literature by studying the optimal payoff choice when the investor assigns asymmetric penalties for gains versus losses with respect to an entire reference distribution – a benchmark – rather than a specific reference point.
To the best of our knowledge, such asymmetry to a benchmark distribution has not yet been explored. To quantify dissimilarities between gains and losses we use the so-called Bregman-Wasserstein (BW) divergence which was introduced in the OT literature by Carlier and Jimenez, (2007) and extensively studied in Kainth et al., (2025). A BW divergence is the minimiser of the Monge-Kantorovich OT problem where the cost function is a Bregman divergence. The BW divergence, which is generated by a convex function $\phi$, is thus an asymmetric generalisation of the (squared) Wasserstein distance. In particular, when $\phi(x)=x^{2}$, it reduces to the squared Wasserstein distance.
We consider a static setting, where admissible payoffs are non-negative measurable functions of a risky asset’s terminal value, i.e., we deal with path-independent payoffs (see also Rüschendorf and Vanduffel, (2020)). Clearly an optimal payoff in the static setting might be surpassed if dynamic trading is permitted, but this outperformance is only apparent if transactions do not bear a cost. In fact, if each transaction has a minimum cost, continuous trading leads to instantaneous bankruptcy and thus is not feasible. Allowing for discrete intermediate trading could mitigate this issue, but it presents major mathematical challenges [^3] Few studies address portfolio maximization under discrete trading with minimum transaction costs. Notable exceptions include Belak et al., (2022) and Bayraktar et al., (2022) who explore the maximization of expected utility in a Black-Scholes market.. The class of path-independent payoffs is extensive in that any distribution function of terminal wealth can be attained and - under technical conditions - the path-independent payoff can be replicated by trading at series of suitable calls and puts [^4] The assumption that all calls and puts are available is arguably as reasonable as the assumption that continuous trading is feasible. To this regard, note that Carr and Chou, (1997) consider this assumption “analogous to the continuous trading assumption permeating the continuous time literature.” (see Breeden and Litzenberger, (1978)).
The paper is structured as follows. Section 2 provides the investors problem description and Section 3 provides our main result, that is the optimal investment strategy. Section 4 discusses examples and Section 5 concludes.
2 Problem Formulation
2.1 Financial Market
There are two assets available in the market, a risk-free asset with current value $B_{0}>0$ and value $B_{T}=B_{0}e^{rT}$ at terminal time $T>0$, where $r>0$ is the risk-free interest rate, and a risky asset with current value $S_{0}>0$ and value at time $T$ that is described by the random variable $S_{T}:\Omega\to[0,\infty)$. We operate within the complete atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$, in which $\mathcal{F}=\sigma(S_{T})$ denotes the $\sigma$-algebra generated by $S_{T}$ and $\mathbb{P}$ is a probability measure, which may be an objective statistical measure derived from historical data or a subjective measure chosen by the investor based on personal beliefs and preferences.
We consider an investor who aims to acquire at $t=0$ a non-negative payoff $X_{T}$ that is a measurable function of the risky asset’s terminal value and lies in $L^{2}$, the space of square-integrable random variables. Furthermore, $X_{T}$ must be affordable given the investor’s initial budget $x_{0}>0$, meaning $X_{T}$ belongs to the set $\mathcal{X}(x_{0})$ given by
$\mathcal{X}(x_{0})\operatorname*{:=}\big\{X_{T}=g(S_{T})~|~g\text{ non-negative, measurable, }{X_{T}\in L^{2},}\,\text{ and }\,c(X_{T})\leq x_{0}\big\}\,$,
where $c(X_{T})$ denotes the cost of $X_{T}$. We assume that a pricing measure $\mathbb{Q}$ exists and that the cost $c(X_{T})$ of $X_{T}$ is given by $c(X_{T}):=\mathbb{E}_{\mathbb{Q}}[e^{-rT}X_{T}]$, where $\mathbb{E}$$_{\mathbb{Q}}$$\cdot$] denotes the expected value under $\mathbb{Q}$}. Equivalently, the cost $c(X_{T})$ can be written as
$c(X_{T})=\mathbb{E}_{\mathbb{P}}[\varphi_{T}X_{T}],$
where $\varphi_{T}:=e^{-rT}\frac{\text{d}\mathbb{Q}}{\text{d}\mathbb{P}}$ is the state price density. For simplicity, we write ${\mathbb{E}}[\cdot]:={\mathbb{E}}_{\mathbb{P}}[\cdot]$, when considering the expected value under $\mathbb{P}$}.
Remark 2.1.
If call options $c_{K}:=\max\{S_{T}-K,0\}$ are traded for all exercise prices $K\geq 0$, then the pricing measure $\mathbb{Q}$ is uniquely determined by these call prices, see e.g., Breeden and Litzenberger, (1978), Ross, (1976), and Nachman, (1988).
In the sequel, all cumulative distribution functions (cdfs) are taken with respect to $\mathbb{P}$. For a given cdf $F$, we denote its (left-continuous) quantile function [^5] The left-continuous quantile function of a cdf $F$ is defined as ${\breve{F}}(u)=\inf\{y\in\mathbb{R}~|~F(y)\geq u\},~0<u<1$. by $\breve{F}$} and for a random variable $Y$, we write $F_{Y}$ and ${\breve{F}}_{Y}$ for its cdf and quantile function, respectively. Furthermore, we assume that the state-price density $\varphi_{T}$ is continuously distributed, satisfies $0<\varphi_{T}<+\infty$ $\mathbb{P}$}-a.s.
As we work with non-negative square-integrable random variables, we define by $\cal{Q}$ the set of quantile functions corresponding to non-negative square-integrable random variables. That is
${\mathcal{Q}}:=\big\{{\breve{G}}:(0,1)\to[0,\infty)\ |\ {\breve{G}}\ \text{is non-decreasing, left-continuous, and }{\int_{0}^{1}\big({\breve{G}}(u)\big)^{2}{\mathrm{d}}u<\infty}\big\}.$
2.2 Expected Utility Preferences
The investor employs a utility function $u(\cdot)$ to attach a value $u(x)$ to each payoff outcome $x$ of $X_{T}$ and evaluates $X_{T}$ by its expected utility $\mathbb{E}[u(X_{T})]$. Throughout, we assume that the utility function $u:[0,\infty)\to\mathbb{R}$ satisfies the following properties:
  1. $i)$
    $u$ is twice continuously differentiable,
  2. ii)ii)
    $u$ is strictly increasing and strictly concave, i. e. $u^{\prime}(x)>0$ and $u^{\prime\prime}(x)<0$ for all $x\in[0,\infty)$, and
  3. $iii)$
    $u$ satisfies the Inada-conditions, i. e.,
    $\lim_{x\to 0}u^{\prime}(x)=\infty\quad\text{and}\quad\lim_{x\to\infty}u^{\prime}(x)=0\,.$
The investor is thus risk averse and moreover the inverse marginal utility function $I:=(u^{\prime})^{-1}:[0,\infty)\to[0,\infty)$ exists and is strictly decreasing and continuously differentiable.
2.3 Bregman-Wasserstein Uncertainty
The investor aims to control the divergence from their payoff to a given benchmark. To formalise this we utilise the notion of Bregman-Wasserstein divergence, for which we first recall the Bregman divergence.
Definition 2.2 (Bregman divergence).
Let $\phi\colon\mathbb{R}\to\mathbb{R}$ be a convex and continuously differentiable function, called a Bregman generator. Then, the Bregman divergence with generator $\phi$ is defined as
$B_{\phi}\big(z_{1},z_{2}\big):=\phi(z_{1})-\phi(z_{2})-\phi^{\prime}(z_{2})(z_{1}-z_{2})\,,\quad z_{1},z_{2}\in\mathbb{R}\,$,
where $\phi^{\prime}(z):=\frac{d}{dz}\phi(z)$ denotes the derivative of $\phi$.
As we work with non-negative random variables, we typically consider Bregman generators $\phi\colon[0,\infty)\to\mathbb{R}$. Note that for the choice $\phi(x)=x^{2}$, it holds that $B_{\phi}(z_{1},z_{2})=(z_{1}-z_{2})^{2}$, i.e., we obtain the squared Euclidean distance. In general, however, the Bregman divergence lacks symmetry (and thus is not a distance). This asymmetry is of great interest in that it makes it possible to penalize positive deviations differently to negative one. Figure 1 illustrates the Bregman divergence for $\phi_{1}(x)=x^{2}$ (blue lines) and $\phi_{2}(x)=x\ln x$, $x>0$ (red lines). By symmetry it holds that $B_{\phi_{1}}(1.5,0.8)=0.49=B_{\phi_{1}}(0.8,1.5)$, while for $\phi_{2}(x)=x\ln x$, we have $B_{\phi_{2}}(1.5,0.8)=0.2429>0.1971=B_{\phi_{2}}(0.8,1.5)$.
Figure 1: The Bregman divergence for the generators $\phi_{1}(x)=x^{2}$ (blue lines) and $\phi_{2}(x)=x\ln x$ (red lines). Left panel displays $B_{\phi_{i}}(1.5,0.8)$ (purple vertical lines) and right panel displays $B_{\phi_{i}}(0.8,1.5)$ (purple vertical lines), $i=1,2$.
With the Bregman divergence, we can define a divergence on the space of cdfs as follows. Let $\Pi(F_{1},F_{2})$ denote the set of all bivariate cdfs with marginal cdfs $F_{1}$ and $F_{2}$, respectively. Pesenti and Vanduffel, (2024) investigate the Monge-Kantorovich optimal transport problem
${\mathscr{B}}[F_{1},F_{2}]:=\inf_{\pi\in\Pi(F_{1},\,F_{2})}\;\left\{\,\int_{\mathbb{R}^{2}}c(z_{1},z_{2})\,\pi({\mathrm{d}}z_{1},{\mathrm{d}}z_{2})\,\right\},$
using as cost function $c(z_{1},z_{2})$ the Bregman divergence $B_{\phi}(z_{1},z_{2})$ and show that the infimum is attained for the comonotonic coupling, i.e., by $\big({\breve{F}}_{1}(U),{\breve{F}}_{2}(U)\big)$, for any $U\sim U(0,1)$. As a consequence, the so-called Bregman-Wasserstein (BW) divergence from $F_{1}$ to $F_{2}$ has representation
${\mathscr{B}}_{\phi}[F_{1},F_{2}]$ $=\int_{0}^{1}B_{\phi}\big({\breve{F}}_{1}(t),{\breve{F}}_{2}(t)\big)\,{\mathrm{d}}t$
$=\int_{0}^{1}\Big(\phi\big({\breve{F}}_{1}(t)\big)-\phi\big({\breve{F}}_{2}(t)\big)-\phi^{\prime}\big({\breve{F}}_{2}(t)\big)\big({\breve{F}}_{1}(t)-{\breve{F}}_{2}(t)\big)\,\Big)\,{\mathrm{d}}t\,$,
which reduces to the 2-Wasserstein distance for $B_{\phi}$ being the squared loss, i.e., when $\phi(x)=x^{2}$. Since the BW divergence only depends on the marginal cdfs, we write ${\mathscr{B}}_{\phi}[F_{1},F_{2}]={\mathscr{B}}_{\phi}({\breve{F}}_{1},{\breve{F}}_{2})$ and use square brackets for cdfs and round brackets for quantile functions.
The asymmetry of the Bregman-Wasserstein divergence reflects how investors evaluate deviations relative to a benchmark. Since the seminal work of Tversky and Kahneman, (1992), it has been well documented that individuals exhibit loss aversion, that is they weigh losses more heavily than gains of equal magnitude. For instance, pension funds or insurance companies prioritize downside protection to ensure solvency, while asset managers may tolerate some underperformance but seek to limit substantial drawdowns relative to benchmarks. The symmetric Wasserstein distance, which penalizes upward and downward deviations equally, cannot capture these asymmetric sensitivities. By contrast, the BW divergence permits a richer class of constraints where negative deviations can be penalized more severely than positive ones, aligning payoff design with observed investor behaviour and regulatory practice.
2.4 Investor’s Optimization Problem
The investor aims to maximize their expected utility under the constraint that their terminal wealth does not diverge too much from the terminal wealth arising from a benchmark (reference) portfolio. Thus, we study the following EUT distance constrained optimization problem:
(P) $\underset{\begin{cases}\ c(X_{T})\leq x_{0}\\[3.00003pt] {\breve{F}}_{X_{T}}\in{\mathcal{Q}}_{\varepsilon}\end{cases}}{\max}{\mathbb{E}}[u(X_{T})]\,$,
where ${\mathcal{Q}}_{\varepsilon}$ is the set of quantile functions that have a BW divergence of at most ${\varepsilon}>0$ to the quantile function of the benchmark’s terminal wealth $\breve{F}_{b}$}, i.e.,
${\mathcal{Q}}_{\varepsilon}:=\Big\{{\breve{G}}~|~{\breve{G}}\in{\mathcal{Q}}\quad\text{and}\quad{\mathscr{B}}_{\phi}({\breve{G}},{\breve{F}_{b}})\leq{\varepsilon}\Big\}\,.$
The value of ${\varepsilon}>0$ is chosen by the investor and we refer to it as the tolerance level. To solve problem (P), we first reduce the set containing all payoffs $X_{T}$ to the subset of payoffs that are anti-monotonic [^6] Two random variables $X$ and $Y$ are called anti-monotonic if for every $\omega_{1}$, $\omega_{2}\in\Omega$ it holds that $(X(\omega_{1})-X(\omega_{2}))(Y(\omega_{1})-Y(\omega_{2}))\leq 0.$ By contrast, they are comonotonic if for every $\omega_{1}$, $\omega_{2}\in\Omega$, $(X(\omega_{1})-X(\omega_{2}))(Y(\omega_{1})-Y(\omega_{2}))\geq 0$. to the state price density $\varphi_{T}$. To see this, assume that some payoff $X_{T}$ distributed with $F$ solves the EUT optimal payoff selection problem (P) [^7] The existence of such a distribution $F$ can be ensured under 2.1 and 3.1; see Theorem 3.5 for a detailed proof. and is not anti-monotonic to the state price density $\varphi_{T}$. We can then obtain a payoff $\hat{X}_{T}$ that has the same cdf $F$ and is strictly cheaper. Indeed, $\hat{X}_{T}:={\breve{F}}(1-F_{\varphi_{T}}(\varphi_{T}))$ has distribution $F$, thus the same expected utility, and satisfies $c(\hat{X}_{T})<c(X_{T})$ by the Hoeffding–Fréchet bounds. This reasoning was first established in Dybvig, (1988) and we refer to Rüschendorf and Vanduffel, (2020) and many of the references therein for more detail. As a consequence, the optimal portfolio selection problem reduces to an optimization problem on real functions under monotonicity constraints (quantile functions).
Theorem 2.3 (Anti-monotonicity of solution).
If the EUT optimization problem (P) has a solution, then there exists a solution $\hat{X}_{T}$ that is anti-monotonic with $\varphi_{T}$.
Payoffs that are anti-monotonic with the state price density $\varphi_{T}$ are called cost-efficient, see e.g., Bernard et al., (2014).
Next, we define the set of quantile functions that have a BW-distance to the reference cdf of at most $\varepsilon$} and whose corresponding cost-efficient payoff costs no more than $x_{0}$. Specifically,
${\mathcal{Q}}_{\varepsilon}(x_{0})\operatorname*{:=}\bigl\{{\breve{G}}\in{\mathcal{Q}}~\big|~{\mathscr{B}}_{\phi}({\breve{G}},{\breve{F}_{b}})\leq{\varepsilon}\;\text{ and }\;{\mathfrak{c}}({\breve{G}})\leq x_{0}\bigr\}\subseteq{\mathcal{Q}}_{\varepsilon}\,$,
where ${\mathfrak{c}}({\breve{G}}):=\int_{0}^{1}{\breve{G}}(t){\breve{F}}_{\varphi_{T}}(1-t)\mathrm{d}t$.
We require the constraint set to be non-empty and the existence of a quantile function with BW divergence strictly smaller than $\varepsilon$}, stated in the following assumption.
Assumption 2.1.
There exists a ${\breve{G}}\in{\mathcal{Q}}_{\varepsilon}(x_{0})$ with ${\mathscr{B}}_{\phi}({\breve{G}},{\breve{F}_{b}})<{\varepsilon}$.
This assumption aligns with the Slater condition (Definition A.4). Given that ${\mathfrak{c}}({\breve{G}})$ is affine in $\breve{G}$}, it is sufficient to assume the existence of a ${\breve{G}}\in{\mathcal{Q}}_{\varepsilon}(x_{0})$ that satisfies the strict BW constraint, see Boyd and Vandenberghe, (2004). This assumption is non-restrictive and is for example fulfilled if the benchmark’s terminal wealth satisfies the budget constraint.
The following result is now classical in the literature on portfolio choice and we omit its proof. [^8] This “quantile reformulation” of the optimal payoff selection problem has been first established in the pioneering work of Dybvig, (1988). It has been further refined and applied in numerous research works including those by Föllmer and Schied, (2004), Carlier and Dana, (2006), Burgert and Rüschendorf, (2006), Jin and Zhou, (2008), Kassberger and Liebmann, (2012), Xu and Zhou, (2013), Bernard et al., 2015b , Von Hammerstein et al., (2014), Xu, (2014), Xu, (2016), Rüschendorf and Vanduffel, (2020), Pesenti and Jaimungal, (2023), He and Jiang, (2021), Wei, (2021), Bi et al., (2023), Magnani et al., (2022), and Bernard et al., 2024a .
Theorem 2.4 (Quantile reformulation).
If $\breve{F}$} is a solution to the optimization problem
($\breve{P}$) $\underset{{\breve{G}}\in{{\mathcal{Q}}_{\varepsilon}(x_{0})}}{\max}\int_{0}^{1}u({\breve{G}}(t))\mathrm{d}t\,$,
then $\hat{X}_{T}={{{\breve{F}}}}(1-F_{\varphi_{T}}(\varphi_{T}))$ solves problem (P).
Solving the EUT optimal payoff choice problem under a BW divergence constraint thus involves addressing an optimization problem concerning functions under monotonicity constraints. In other words, it entails finding a suitable quantile function.
3 Optimal Payoff
In this section we solve the EUT optimisation problem ($\breve{P}$). For this we first establish that problem ($\breve{P}$) is well-posed and that there exists a unique solution.
3.1 Existence and Uniqueness
To obtain existence and uniqueness of the optimal payoff, we first establish that problem ($\breve{P}$) is well-posed, i.e., its value is finite. Before proving the well-posedness of problem ($\breve{P}$), we introduce the following definition.
Definition 3.1 ($\vartheta$-strongly convex).
A function $f:\mathbb{R}\to\mathbb{R}$ is $\vartheta$-strongly convex if there exist $\vartheta>0$ such that
(1) $f(z_{1})\geq f(z_{2})+f^{\prime}(z_{2})(z_{1}-z_{2})+\frac{\vartheta}{2}(z_{1}-z_{2})^{2}\,,\quad z_{1},z_{2}\in\mathbb{R}\,.$
For any Bregman generator $\phi$ that is $\vartheta$-strongly convex, the induced divergence is naturally linked to the squared Euclidean distance and to the Wasserstein divergence. Specifically, recalling Definition 2.2, we have
$B_{\phi}\big(z_{1},z_{2}\big)\geq\frac{\vartheta}{2}(z_{1}-z_{2})^{2}=\frac{\vartheta}{2}B_{\phi_{1}}\big(z_{1},z_{2}\big)\,,\quad z_{1},z_{2}\in\mathbb{R}\,$,
where $\phi_{1}(x)=x^{2}$. This property plays a key role in establishing integrability and thus well-posedness in our framework. The next assumption provides three different cases, under which the problem ($\breve{P}$) is well-posed.
Assumption 3.1.
Let at least one of the following conditions hold:
  1. $i)$
    The state-price density satisfies $\operatorname*{ess\,inf}\varphi_{T}>0$.
  2. ii)ii)
    The Bregman generator $\phi$ is $\vartheta$-strongly convex and the benchmark belongs to $L^{2}$.
  3. $iii)$
    The Bregman generator $\phi$ is strictly convex and the benchmark is bounded.
Proposition 3.2 (Well-posedness).
Under 3.1 problem ($\breve{P}$) is well-posed.
Proof 3.3.
For any feasible solution $\breve{F}$ of problem ($\breve{P}$), we have by Jensen’s inequality
$\int_{0}^{1}u(\breve{F}(t))\mathrm{d}t\leq u\Big(\int_{0}^{1}\breve{F}(t)\mathrm{d}t\Big)\,.$
This implies that if $\int_{0}^{1}\breve{F}(t)\,\mathrm{d}t<\infty$, then the value of problem ($\breve{P}$), denoted by $V$, is finite. For case $i)$, we have
$\int_{0}^{1}\breve{F}(t)\mathrm{d}t\leq\frac{1}{\displaystyle\lim_{t\to 1}\breve{F}_{\varphi_{T}}(1-t)}\int_{0}^{1}\breve{F}(t)\breve{F}_{\varphi_{T}}(1-t)\mathrm{d}t\leq\frac{x_{0}}{\displaystyle\lim_{t\to 1}\breve{F}_{\varphi_{T}}(1-t)}<\infty\,.$
For case ii)ii), we note that as $\phi$ is $\vartheta$-strong convex, we have $B_{\phi}\big(z_{1},z_{2}\big)\geq\frac{\vartheta}{2}(z_{1}-z_{2})^{2}$ as well as ${\mathscr{B}}_{\phi}(\breve{F}_{1},\breve{F}_{2})\geq\frac{\vartheta}{2}{\mathscr{B}}_{\phi_{1}}(\breve{F}_{1},\breve{F}_{2})$, with $\phi_{1}(x)=x^{2}$. Then, from the BW constraint, we obtain
$\frac{\vartheta}{2}{\mathscr{B}}_{\phi_{1}}(\breve{F},\breve{F}_{b})\leq{\mathscr{B}}_{\phi}(\breve{F},\breve{F}_{b})\leq\varepsilon\,$,
that is
${\mathscr{B}}_{\phi_{1}}(\breve{F},\breve{F}_{b})=\int_{0}^{1}\big(\breve{F}(t)-\breve{F}_{b}(t)\big)^{2}\mathrm{d}t\leq\frac{2\varepsilon}{\vartheta}.$
Moreover, we have
$\int_{0}^{1}\breve{F}(t)\mathrm{d}t-\int_{0}^{1}\breve{F}_{b}(t)\mathrm{d}t$ $\leq\Big|\int_{0}^{1}\breve{F}(t)\mathrm{d}t-\int_{0}^{1}\breve{F}_{b}(t)\mathrm{d}t\Big|$
$\leq\sqrt{\int_{0}^{1}\big(\breve{F}(t)-\breve{F}_{b}(t)\big)^{2}\mathrm{d}t}$
$\leq\sqrt{\frac{2\varepsilon}{\vartheta}}\,$,
where the second inequation follows from the Cauchy-Schwarz inequality (applied to the functions $\breve{F}(\cdot)-\breve{F}_{b}(\cdot)$ and $1$), and we obtain
$\int_{0}^{1}\breve{F}(t)\mathrm{d}t\leq\int_{0}^{1}\breve{F}_{b}(t)\mathrm{d}t+\sqrt{\frac{2\varepsilon}{\vartheta}}<\infty.$
For case $iii)$, let the Bregman generator $\phi$ be strictly convex and the benchmark be bounded. Since the benchmark is bounded, then for all $t\in(0,1)$, we have $m_{b}\leq\breve{F}_{b}(t)\leq M_{b}$, where $m_{b}$ and $M_{b}$ are two given constants. Consider $x>0$ and $y=\breve{F}_{b}(t)$, from the definition of Bregman divergence, we have
$B_{\phi}\big(x,y\big)=$ $\ \phi(x)-\phi(y)-\phi^{\prime}(y)(x-y)$
$=$ $\ \phi(x)-\phi^{\prime}(y)x+\big(y\phi^{\prime}(y)-\phi(y)\big),$
where the terms $\phi^{\prime}(y)$ and $y\phi^{\prime}(y)-\phi(y)$ are bounded because $y\in[m_{b},M_{b}]$ and $\phi,\phi^{\prime}$ are continuous on the compact set.
Let $C_{1}=\phi^{\prime}(M_{b})$ and $C_{2}=\sup_{y\in[m_{b},M_{b}]}|y\phi^{\prime}(y)-\phi(y)|$, we obtain
$B_{\phi}\big(x,y\big)\geq\phi(x)-C_{1}x-C_{2}.$
Note that $\phi$ is strictly convex, its derivative $\phi^{\prime}$ is strictly increasing. Hence,
$\lim_{x\to\infty}\phi^{\prime}(x)=L\in(C_{1},\infty].$
Next, choose $\delta\in(0,L-C_{1})$ (if $L=\infty$, take any $\delta\in(0,\infty)$). Then there exists $a_{0}>M_{b}$ such that $\phi^{\prime}(x)\geq C_{1}+\delta$ for all $x\geq a_{0}$. By the convexity of $\phi(x)$, we have
$\phi(x)\geq\phi(a_{0})+\phi^{\prime}(a_{0})(x-a_{0})\geq\phi(a_{0})+(C_{1}+\delta)(x-a_{0})\,,\quad x\geq a_{0}.$
Therefore,
$\phi(x)-C_{1}x-C_{2}\geq$ $\ \phi(a_{0})+(C_{1}+\delta)(x-a_{0})-C_{1}x-C_{2}$
$=$ $\ \delta x-((C_{1}+\delta)a_{0}-\phi(a_{0}))-C_{2}\,,\quad x\geq a_{0}.$
Taking $k_{0}=\delta$ and $C_{0}=\max\{C_{2}+(C_{1}+\delta)a_{0}-\phi(a_{0}),0\}$, we obtain that for all $x>a_{0}$ and all $y\in[m_{b},M_{b}]$,
$B_{\phi}\big(x,y\big)\geq k_{0}x-C_{0}.$
Let $\mathcal{S}=\{t\in(0,1)\ |\ \breve{F}(t)>a_{0}\}$ and $\overline{\mathcal{S}}=\{t\in(0,1)\ |\ \breve{F}(t)\leq a_{0}\}$, consider the BW constraint,
$k_{0}\int_{\mathcal{S}}\breve{F}(t)\mathrm{d}t-C_{0}$ $\leq\int_{\mathcal{S}}\big(k_{0}\breve{F}(t)-C_{0}\big)\mathrm{d}t$
$\leq\int_{\mathcal{S}}B_{\phi}\big(\breve{F}(t),\breve{F}_{b}(t)\big)\mathrm{d}t$
$\leq{\mathscr{B}}_{\phi}(\breve{F},\breve{F}_{b})$
$\leq\varepsilon,$
where the first inequality arises from the measure of $\mathcal{S}$ being at most 1. Thus we obtain
$\int_{\mathcal{S}}\breve{F}(t)\mathrm{d}t\leq\frac{C_{0}+\varepsilon}{k_{0}}\qquad\text{and}\qquad\int_{\overline{\mathcal{S}}}\breve{F}(t)\mathrm{d}t\leq a_{0}.$
Thus, we derive that
$\int_{0}^{1}\breve{F}(t)\mathrm{d}t=\int_{\mathcal{S}}\breve{F}(t)\mathrm{d}t+\int_{\overline{\mathcal{S}}}\breve{F}(t)\mathrm{d}t\leq\frac{C_{0}+\varepsilon}{k_{0}}+a_{0}<\infty.$
In conclusion, we showed that in all three cases, any feasible solution has a finite expected utility. Thus problem ($\breve{P}$) is well-posed.
Remark 3.4.
In practice, if a Bregman generator $\phi$ does not satisfy strong convexity—for example, $\phi_{2}(x)=x\ln x$ or the threshold-based generator introduced in Definition 4.3 —strong convexity can be enforced by adding a regularization term of the form $\frac{\vartheta}{2}x^{2}$, where $\vartheta>0$ is sufficiently small (e.g., $10^{-8}$), to the Bregman generator. For instance, for $\phi_{2}(x)$, one may consider the regularized version $\phi_{2,\vartheta}(x)=x\ln x+\frac{\vartheta}{2}x^{2}$. Conceptually, this regularization adds a weighted Wasserstein-type constraint, $\frac{\vartheta}{2}{\mathscr{B}}_{\phi_{1}}({\breve{F}}_{1},{\breve{F}}_{2})$ with $\phi_{1}(x)=x^{2}$, to the original Bregman–Wasserstein constraint ${\mathscr{B}}_{\phi_{2}}({\breve{F}}_{1},{\breve{F}}_{2})$, i.e.,
(2) ${\mathscr{B}}_{\phi_{2,\vartheta}}({\breve{F}}_{1},{\breve{F}}_{2})={\mathscr{B}}_{\phi_{2}}({\breve{F}}_{1},{\breve{F}}_{2})+\frac{\vartheta}{2}{\mathscr{B}}_{\phi_{1}}({\breve{F}}_{1},{\breve{F}}_{2}),$
thereby ensuring the well-posedness of the problem. As observed from (2), when $\vartheta$ is sufficiently small, the deviation penalty is primarily governed by ${\mathscr{B}}_{\phi_{2}}({\breve{F}}_{1},{\breve{F}}_{2})$.
Theorem 3.5 (Existence and uniqueness).
Let Assumptions 2.1 and 3.1 be satisfied. Then there exists a unique solution, denoted by $\breve{F}$$^{*}$, of Problem ($\breve{P}$). Moreover, at least one of the constraints, i.e., either the budget constraint or the BW constraint, is binding for $\breve{F}$$^{*}$.
Proof 3.6.
The proof is broken into three parts. First, we show that the constraint set ${\mathcal{Q}}_{\varepsilon}(x_{0})$ is convex. Second, we prove existence and uniqueness of the solution, and finally, we show that at least one constraint is binding.
Part 1. Convexity of constraint set. Take ${\breve{G}}_{1},{\breve{G}}_{2}\in{\mathcal{Q}}_{\varepsilon}(x_{0})$ and define the convex combination ${\breve{G}}_{\omega}:=\omega{\breve{G}}_{1}+(1-\omega){\breve{G}}_{2}$ for $\omega\in[0,1]$. Clearly, $\int_{0}^{1}{\breve{G}}_{\omega}(t){\breve{F}}_{\varphi_{T}}(1-t)\mathrm{d}t\leq x_{0}$. Furthermore, as the Bregman divergence is convex in its first argument, we have
${\mathscr{B}}_{\phi}({\breve{G}}_{\omega},{\breve{F}_{b}})$ $=\int_{0}^{1}B_{\phi}\big({\breve{G}}_{\omega}(t),{\breve{F}_{b}}(t)\big)\,{\mathrm{d}}t$
$\leq\omega\int_{0}^{1}B_{\phi}\big({\breve{G}}_{1}(t),{\breve{F}_{b}}(t)\big)\,{\mathrm{d}}t+(1-\omega)\int_{0}^{1}B_{\phi}\big({\breve{G}}_{2}(t),\ {\breve{F}_{b}}(t)\big)\,{\mathrm{d}}t$
$=\omega\,{\mathscr{B}}_{\phi}({\breve{G}}_{1},{\breve{F}_{b}})+(1-\omega)\,{\mathscr{B}}_{\phi}({\breve{G}}_{2},{\breve{F}_{b}})$
$\leq{\varepsilon}\,$,
where the last inequality follows since ${\breve{G}}_{1},{\breve{G}}_{2}\in{\mathcal{Q}}_{\varepsilon}(x_{0})$. Hence ${\breve{G}}_{\omega}\in{\mathcal{Q}}_{\varepsilon}(x_{0})$ and ${\mathcal{Q}}_{\varepsilon}(x_{0})$ is a convex set.
Part 2. Existence and uniqueness. Let $V$ denote the value of problem ($\breve{P}$). Since $V<\infty$ (3.1) and the feasible region is nonempty (2.1), there exist a sequence $(\breve{G}_{n})_{n\geq 1}\subseteq\mathcal{Q}_{\varepsilon}(x_{0})$ such that
(3) $\int_{0}^{1}u(\breve{G}_{n}(t))\mathrm{d}t>V-\frac{1}{n},\quad\forall n\geq 1.$
Theorem 3.7 in Balder, (1984) states that if $(\breve{G}_{n})_{n\geq 1}$ satisfies a generalised tightness condition and $(u\circ\breve{G}_{n})_{n\geq 1}$ is uniformly integrable, then there exists a convergent subsequence. This means that there exists ${\breve{G}}^{*}\in\mathcal{Q}_{\varepsilon}(x_{0})$ to which a subsequence converges pointwise.
The generalised tightness condition is as follows: The sequence $(\breve{G}_{n})_{n\geq 1}$ is tight if there exists a non-negative normal integrand $l:(0,1)\times[0,\infty)\to[0,\infty)$ such that the following hold:
(1) For every $t\in(0,1)$, $\beta\in\mathbb{R}$, the set $\{s\in[0,\infty):l(t,s)\leq\beta\}$ is compact.
(2) $\sup_{n\geq 1}\int_{0}^{1}l(t,\breve{G}_{n}(t))\mathrm{d}t<\infty$.
Defining $l(t,s)=\breve{F}_{\varphi_{T}}(1-t)s\geq 0$, we obtain that (for fixed $t$ and $\beta$) the set is $\{s\in[0,\infty)\ |\ l(t,s)\leq\beta\}=[0,\frac{\beta}{\breve{F}_{\varphi_{T}}(1-t)}]$ indeed a compact set in $0,\infty)$. Moreover,
$\int_{0}^{1}l(t,\breve{G}_{n}(t))\mathrm{d}t=\int_{0}^{1}\breve{G}_{n}(t)\breve{F}_{\varphi_{T}}(1-t)\mathrm{d}t\leq x_{0}<\infty\,$,
showing that $(\breve{G}_{n})_{n\geq 1}$ is tight.
Next, we prove that $(u\circ\breve{G}_{n})_{n\geq 1}$ is uniformly integrable. According to the assumption of the utility function $u$ (see the beginning of Section 2.2), we have $\lim_{x\to\infty}\frac{u^{-1}(x)}{x}=\infty$. By de la Vallée Poussin criterion with the convex function $u^{-1}$, if we show that
$\int_{0}^{1}u^{-1}(u(\breve{G}_{n}(t)))\mathrm{d}t=\int_{0}^{1}\breve{G}_{n}(t)\mathrm{d}t,$
is uniformly bounded, then the sequence $(u\circ\breve{G}_{n})_{n\geq 1}$ is uniformly integrable. The proof that the ${\breve{G}}_{n}$, $n\geq 1$, have finite expectation follows from 3.1. Indeed in the proof of Proposition 3.2, we showed that for each case in 3.1, the expectation of any feasible solution is uniformly bounded from above. Consequently, there exists a constant $C$ such that, $\int_{0}^{1}{\breve{G}}_{n}(t)\,\mathrm{d}t<C<\infty$ for all cases of 3.1. Hence, the sequence $(u\circ\breve{G}_{n})_{n\geq 1}$ is uniformly integrable in all three cases.
Thus, we have shown that for each of the three cases in 3.1 the conditions of Theorem 3.7 in Balder, (1984) are satisfied. Thus there exist a subsequence $(\breve{G}_{n_{k}})_{k\geq 1}$ that converges to $\breve{G}^{*}$ pointwise, and we have
$\int_{0}^{1}u(\breve{G}^{*}(t))\mathrm{d}t>V-\frac{1}{n_{k}},$
for all $k$. If $\breve{G}^{*}$ is feasible, then $\int_{0}^{1}u(\breve{G}^{*}(t))\mathrm{d}t=V$ and $\breve{G}^{*}$ attains the maximum value of problem ($\breve{P}$). It remains to verify that $\breve{G}^{*}$ is feasible for problem ($\breve{P}$). This follows immediately from Fatou’s Lemma, since
$\int_{0}^{1}\breve{G}^{*}(t)\breve{F}_{\varphi_{T}}(1-t)\mathrm{d}t\leq\liminf_{k\to\infty}\int_{0}^{1}\breve{G}_{n_{k}}(t)\breve{F}_{\varphi_{T}}(1-t)\mathrm{d}t\leq x_{0},$
$\int_{0}^{1}B_{\phi}\big({\breve{G}}^{*}(t),{\breve{F}_{b}}(t)\big){\mathrm{d}}t\leq\liminf_{k\to\infty}\int_{0}^{1}B_{\phi}\big({\breve{G}}_{n_{k}}(t),{\breve{F}_{b}}(t)\big){\mathrm{d}}t\leq\varepsilon.$
Finally, uniqueness follows from the strict concavity of the utility function.
Part 3. At least one of the constraints is binding. Assume that the optimal quantile function $\breve{F}$} is in the interior of ${\mathcal{Q}}_{\varepsilon}(x_{0})$, i.e., ${\mathscr{B}}_{\phi}({\breve{F}},{\breve{F}_{b}})={\varepsilon}^{\prime}<{\varepsilon}$ and ${\mathfrak{c}}({\breve{F}})=x_{0}^{\prime}<x_{0}$. Next, define the quantile function ${\breve{F}}_{c}(t):={\breve{F}}(t)+c$ for $c>0$, then we have
${\mathscr{B}}_{\phi}({\breve{F}}_{c},{\breve{F}_{b}})=$ $\int_{0}^{1}\Big(\phi\big({\breve{F}}(t)+c\big)-\phi\big({\breve{F}_{b}}(t)\big)-\phi^{\prime}\big({\breve{F}_{b}}(t)\big)\big({\breve{F}}(t)+c-{\breve{F}_{b}}(t)\big)\ \Big)\mathrm{d}t$
$=$ ${\mathscr{B}}_{\phi}({\breve{F}},{\breve{F}_{b}})+\int_{0}^{1}\left(\phi\big({\breve{F}}(t)+c\big)-\phi\big({\breve{F}}(t)\big)-c\,\phi^{\prime}\big({\breve{F}_{b}}(t)\big)\right)\mathrm{d}t$
$=$ ${\varepsilon}^{\prime}+\int_{0}^{1}\left(\phi\big({\breve{F}}(t)+c\big)-\phi\big({\breve{F}}(t)\big)-c\,\phi^{\prime}\big({\breve{F}_{b}}(t)\big)\right)\mathrm{d}t,$
and
${\mathfrak{c}}({\breve{F}}_{c})={\mathfrak{c}}({\breve{F}})+c\int_{0}^{1}{\breve{F}}_{\varphi_{T}}(1-t)\mathrm{d}t=x_{0}^{\prime}+c\int_{0}^{1}{\breve{F}}_{\varphi_{T}}(1-t)\mathrm{d}t\,.$
Note that we have $\phi\big({\breve{F}}(t)+c\big)-\phi\big({\breve{F}}(t)\big)-c\,\phi^{\prime}\big({\breve{F}_{b}}(t)\big)\geq 0$, by convexity of $\phi(\cdot)$, and that $\int_{0}^{1}{\breve{F}}_{\varphi_{T}}(1-t)\mathrm{d}t>0$. Furthermore
$\lim_{c\to 0}\int_{0}^{1}\left(\phi\big({\breve{F}}(t)+c\big)-\phi\big({\breve{F}}(t)\big)-c\,\phi^{\prime}\big({\breve{F}_{b}}(t)\big)\right)\mathrm{d}t=\lim_{c\to 0}c\int_{0}^{1}{\breve{F}}_{\varphi_{T}}(1-t)\mathrm{d}t=0,$
Since for each $c>0$ the above integrals are non-negative and by their continuity with respect to $c$, there exists a $\tilde{c}>0$ such that ${\varepsilon}^{\prime}<{\mathscr{B}}_{\phi}({\breve{F}}_{\tilde{c}},{\breve{F}_{b}})\leq{\varepsilon}$ and $x_{0}^{\prime}<{\mathfrak{c}}({\breve{F}}_{\tilde{c}})\leq x_{0}$, which implies that ${\breve{F}}_{\tilde{c}}(t)\in{\mathcal{Q}}_{\varepsilon}(x_{0})$. Moreover, it holds that $\int_{0}^{1}u({\breve{F}}_{\tilde{c}}(t))\mathrm{d}t>\int_{0}^{1}u({\breve{F}}(t))\mathrm{d}t$, which contradicts optimality of $\breve{F}$}. Therefore, at least one of the constraints is binding.
3.1 provides three cases, when the optimisation problem ($\breve{P}$) is well-posed. Case 1 relies solely on assumption concerning the state price density $\varphi_{T}$, and can be viewed as the counterpart, under expected utility, of the portfolio selection framework in Ghossoub and Zhu, (2025). Cases 2 and 3 rely on the BW distance constraint, which effectively restricts the feasible portfolio set to a neighborhood around the benchmark.
3.2 Optimal Quantile Function
Recall that the set $\mathcal{Q}$ consists of non-negative and square-integrable quantile functions. Thus, we can view $\mathcal{Q}$ as a convex subset of $L^{2}((0,1))$. Endowed with the $L^{2}$-norm, $L^{2}((0,1))$ forms a Banach space, which allows us to apply the theory of convex optimisation to problem ($\breve{P}$), provided that the objective function and the constraints meet suitable regularity conditions. A concise review of the relevant results in convex optimisation is provided in Appendix A.
In this section, we work under Assumptions 2.1 and 3.1 and rewrite problem ($\breve{P}$) as follows:
(4) $\inf_{{\breve{G}}\in\mathcal{Q}}\left\{\int_{0}^{1}-u\big({\breve{G}}(t)\big){\mathrm{d}}t\ \Big|\ {\mathfrak{c}}({\breve{G}})-x_{0}\leq 0,\ {\mathscr{B}}_{\phi}({\breve{G}},{\breve{F}_{b}})-{\varepsilon}\leq 0\right\}.$
In the proof of Theorem 3.5 (Part 1.), we establish the convexity of the distance constraint. Moreover, the objective function is convex and the budget constraint linear. For the problem (4) to satisfies the conditions in Theorem A.1 and Theorem A.5, we need to define the objective and constraint functions on the entire space $L^{2}((0,1))$. For this, let ${,+}\bar{\mathbb{R}}:=\mathbb{R}\cup\{$-$\infty$$\infty\},$ and define the functional $\bar{\mathfrak{u}}:L^{2}((0,1))\to\bar{\mathbb{R}}$ by
$\bar{\mathfrak{u}}({\breve{G}}):=\begin{cases}-\int_{0}^{1}u({\breve{G}}(t))\mathrm{d}t,\qquad&\text{if}\quad{\breve{G}}\in\mathcal{Q},\\[5.0pt] +\infty&\text{if}\quad{\breve{G}}\in L^{2}((0,1))/\mathcal{Q}\end{cases}$
and similarly
$\bar{{\mathfrak{c}}}({\breve{G}}):$ $=\begin{cases}{\mathfrak{c}}({\breve{G}}),\qquad&\text{if}\quad{\breve{G}}\in\mathcal{Q},\\[5.0pt] +\infty&\text{if}\quad{\breve{G}}\in L^{2}((0,1))/\mathcal{Q}\end{cases}\qquad\text{and}\qquad$
$\bar{{\mathscr{B}}}({\breve{G}},{\breve{F}_{b}}):$ $=\begin{cases}{\mathscr{B}}({\breve{G}},{\breve{F}_{b}}),\qquad&\text{if}\quad{\breve{G}}\in\mathcal{Q},\\[5.0pt] +\infty&\text{if}\quad{\breve{G}}\in L^{2}((0,1))/\mathcal{Q}\,.\end{cases}$
Thus problem (4) is equivalent to
(5) $\operatorname*{arg\,min}_{{\breve{G}}\in L^{2}((0,1))}\bigg\{\bar{\mathfrak{u}}({\breve{G}})+\lambda\,\bar{{\mathfrak{c}}}({\breve{G}})+\mu\,\bar{{\mathscr{B}}}_{\phi}({\breve{G}},{\breve{F}_{b}})\bigg\},$
where we optimise over ${\mathcal{L}}^{2}((0,1))$. Now, we can apply Theorem A.1 to obtain a solution and if the solution belongs to $\mathcal{Q}$, it is also optimal for problem (9). For ease of representation, we simply write $\int_{0}^{1}u({\breve{G}}(t)){\mathrm{d}}t$, ${\mathfrak{c}}({\breve{G}})$, and ${\mathscr{B}}({\breve{G}},{\breve{F}_{b}})$.
As a direct consequence of 2.1, problem (4) satisfies the Slater condition (see Definition A.4) and the regularity properties outlined in Theorem A.1. Thus, by Theorem A.5 the existence of a solution is equivalent to the existence of a pair of real numbers $(\lambda^{*},\mu^{*})$ such that the KKT conditions are satisfied. Hence, if ${\breve{G}}^{*}{{\in\mathcal{Q}}}$ is a solution to problem (4), Theorem A.1 ensures the existence of $\lambda^{*},\mu^{*}\geq 0$ such that
(6) $-\int_{0}^{1}u({\breve{G}}^{*}(t)){\mathrm{d}}t$ $\leq-\int_{0}^{1}u({\breve{G}}(t)){\mathrm{d}}t+\lambda^{*}\left({\mathfrak{c}}({\breve{G}})-x_{0}\right)+\mu^{*}\left({\mathscr{B}}_{\phi}({\breve{G}},{\breve{F}_{b}})-{\varepsilon}\right)$
for all ${\breve{G}}\in\cal{Q}$. Additionally, by Theorem A.1, statement 3., we have
(7) $\lambda^{*}\left({\mathfrak{c}}\big({\breve{G}}^{*}\big)-x_{0}\right)=0\,\quad\text{and}\quad\mu^{*}\left({\mathscr{B}}_{\phi}\big({\breve{G}}^{*},{\breve{F}}_{b}\big)-{\varepsilon}\right)=0\,.$
Adding these terms to the left-hand side of (6) and cancelling the $x_{0}$ and $\varepsilon$} terms yields:
$\int_{0}^{1}-u({\breve{G}}^{*}(t)){\mathrm{d}}t+\lambda^{*}{\mathfrak{c}}({\breve{G}}^{*})+\mu^{*}{\mathscr{B}}_{\phi}({\breve{G}}^{*},{\breve{F}_{b}})\leq\int_{0}^{1}-u({\breve{G}}(t)){\mathrm{d}}t+\lambda^{*}{\mathfrak{c}}({\breve{G}})+\mu^{*}{\mathscr{B}}_{\phi}({\breve{G}},{\breve{F}_{b}})\,$
for all ${\breve{G}}\in\mathcal{Q}$. Therefore, $\breve{G}$$^{*}$ must also be a solution to the problem
(8) $\operatorname*{arg\,min}_{{\breve{G}}\in\mathcal{Q}}\bigg\{$ $\int_{0}^{1}-u({\breve{G}}(t)){\mathrm{d}}t+\lambda^{*}\,{\mathfrak{c}}({\breve{G}})+\mu^{*}\,{\mathscr{B}}_{\phi}({\breve{G}},{\breve{F}_{b}})\bigg\}\,.$
Next, Theorem A.5 shows that any solution of (8) that satisfies the slackness conditions (7), is also optimal for problem (6). Thus, for arbitrary given multipliers $\lambda,\mu\geq 0$, we next solve
$\operatorname*{arg\,min}_{{\breve{G}}\in\mathcal{Q}}$ $\bigg\{\int_{0}^{1}-u({\breve{G}}(t)){\mathrm{d}}t+\lambda\,{\mathfrak{c}}({\breve{G}})+\mu\,{\mathscr{B}}_{\phi}({\breve{G}},{\breve{F}_{b}})\bigg\}$
(9) $=\operatorname*{arg\,min}_{{\breve{G}}\in\mathcal{Q}}\bigg\{\int_{0}^{1}-u({\breve{G}}(t))+\lambda{\breve{G}}(t){\breve{F}}_{\varphi_{T}}(1-t)+\mu\big(\phi\big({\breve{G}}(t)\big)-\phi^{\prime}\big({\breve{F}_{b}}(t)\big){\breve{G}}(t)\big){\mathrm{d}}t\bigg\}\,.$
Fix a value of $t\in(0,1)$, and consider the integrand of problem (9) as a function of $y$, defined as
(10) $h_{t}(y):=\ m(y)+\lambda y{\breve{F}}_{\varphi_{T}}(1-t)-\mu\phi^{\prime}\big({\breve{F}_{b}}(t)\big)y\,$,
where we set $m(y):=-u(y)+\mu\phi(y)$. Then, we consider the pointwise optimiser of $h_{t}(\cdot)$, that is for each $t\in(0,1)$ we solve
(11) $\operatorname*{arg\,min}_{y\in[0,\infty]}\;h_{t}(y)\,.$
As a solution to (11) minimizes the integral in (9) pointwise, it is a potential solution to problem (9), if the solution to (11) is a quantile function, i.e., non-decreasing and left-continuous. In the following, we first solve (11) and second verify that its solution is indeed a quantile function.
Proposition 3.7.
For $\lambda,\mu\geq 0$ and for all $t\in(0,1)$, let
(12) ${\breve{G}}_{\lambda,\mu}(t):$ $=\inf\left\{y\geq 0\ |\ \tfrac{\partial}{\partial y}\,h_{t}(y)\geq 0\right\}\,$,
where by convention we set $\inf\emptyset:=+\infty$. Then
(13) ${\breve{G}}_{\lambda,\mu}=\operatorname*{arg\,min}_{H\colon(0,1)\to[0,\infty]}\int_{0}^{1}-u\big(H(t)\big){\mathrm{d}}t+\lambda\,{\mathfrak{c}}(H)+\mu\,{\mathscr{B}}_{\phi}(H,{\breve{F}_{b}})\,.$
Proof 3.8.
[…]
(14) $\lim_{y\to 0}\tfrac{\partial}{\partial y}\,h_{t}(y)=\lim_{y\to 0}-u^{\prime}(y)+\lambda{\breve{F}}_{\varphi_{T}}(1-t)+\mu\left(\phi^{\prime}(y)-\phi^{\prime}\big({\breve{F}_{b}}(t)\big)\right)<0\,$,
since by convexity of $\phi$, it holds that $\lim_{y\to 0}\phi^{\prime}(y)<+\infty$. Indeed if $\lim_{y\to 0}\phi^{\prime}(y)=+\infty$, then $\phi^{\prime}(y)=+\infty$ for all $y\geq 0$, since $\phi^{\prime}$ is non-decreasing. From the above inequality, we obtain that for each $t\in(0,1)$, the integrand $h_{t}(y)$ is decreasing at $y=0$, which implies that the argmin is strictly greater than zero for all $t\in(0,1)$.
Next, we consider the case when $\lim_{y\to+\infty}\tfrac{\partial}{\partial y}\,h_{t}(y)\leq 0$, that is $\tfrac{\partial}{\partial y}\,h_{t}(y)$ does not change sign and remains non-positive. Consequently, $\tfrac{\partial}{\partial y}\,h_{t}(y)<0$ for all $y\in[0,\infty)$, implying that $h_{t}(y)$ is strictly decreasing on $0,\infty)$ and attains its minimum at $\infty$. Hence, we conclude $h_{t}(y)$ is minimized by (12).
The next proposition establishes that ${\breve{G}}_{\lambda,\mu}$ is a quantile function.
Proposition 3.9.
For all $\lambda,\mu\geq 0$, ${\breve{G}}_{\lambda,\mu}(\cdot)$ is non-decreasing and left-continuous.
Proof 3.10.
We first show non-decreasingness and then left-continuity.
Part 1. Non-decreasingness. First, we show that if there exists a $t_{1}\in(0,1)$ such that ${\breve{G}}_{\lambda,\mu}(t_{1})=+\infty$, then ${\breve{G}}_{\lambda,\mu}(t_{2})=+\infty$ for all $t_{2}\geq t_{1}$. From the proof of Proposition 3.7, we know that ${\breve{G}}_{\lambda,\mu}(t_{1})=+\infty$ is equivalent to
$\frac{\partial}{\partial y}h_{t_{1}}(y)=m^{\prime}(y)+\lambda{\breve{F}}_{\varphi_{T}}(1-t_{1})-\mu\phi^{\prime}\big({\breve{F}_{b}}(t_{1})\big)<0,$
for all $y\in[0,\infty)$. Since ${\breve{F}}_{\varphi_{T}}(1-t)$ is decreasing and $\phi^{\prime}\big({\breve{F}_{b}}(t)\big)$ is non-decreasing with respect to $t$, we also obtain that
$\frac{\partial}{\partial y}h_{t_{2}}(y)=m^{\prime}(y)+\lambda{\breve{F}}_{\varphi_{T}}(1-t_{2})-\mu\phi^{\prime}\big({\breve{F}_{b}}(t_{2})\big)<0,$
for all $y\in[0,\infty)$, and thus ${\breve{G}}_{\lambda,\mu}(t_{2})=+\infty$. Thus, we conclude that ${\breve{G}}_{\lambda,\mu}(t_{1})=+\infty$, implies that ${\breve{G}}_{\lambda,\mu}(t_{2})=+\infty$ for all $t_{2}\geq t_{1}$.
Next, we verify that ${\breve{G}}_{\lambda,\mu}(t)$ is non-decreasing within the set $\{t\in(0,1)\ |\ {\breve{G}}_{\lambda,\mu}(t)<+\infty\}$. For $t_{1}<t_{2}$ in that set, we have by continuity of $\frac{\partial}{\partial y}h_{t}(y)$ that
$\tfrac{\partial}{\partial y}h_{t_{1}}(y)=m^{\prime}({\breve{G}}_{\lambda,\mu}(t_{1}))+\lambda{\breve{F}}_{\varphi_{T}}(1-t_{1})-\mu\phi^{\prime}\big({\breve{F}_{b}}(t_{1})\big)=0\,,\quad\text{and}$
$\tfrac{\partial}{\partial y}h_{t_{2}}(y)=m^{\prime}({\breve{G}}_{\lambda,\mu}(t_{2}))+\lambda{\breve{F}}_{\varphi_{T}}(1-t_{2})-\mu\phi^{\prime}\big({\breve{F}_{b}}(t_{2})\big)=0.$
Subtracting the two equations and rearranging, we have
(15) $\begin{split}m^{\prime}({\breve{G}}_{\lambda,\mu}(t_{2}))-m^{\prime}({\breve{G}}_{\lambda,\mu}(t_{1}))=-\lambda\left({\breve{F}}_{\varphi_{T}}(1-t_{2})-{\breve{F}}_{\varphi_{T}}(1-t_{1})\right)\\ +\mu\left(\phi^{\prime}\big({\breve{F}_{b}}(t_{2})\big)-\phi^{\prime}\big({\breve{F}_{b}}(t_{1})\big)\right)\,.\end{split}$
Note that the right hand side of (15) is non-negative by the monotonicity of the quantile functions ${\breve{F}}_{\varphi_{T}}(\cdot)$ and ${\breve{F}_{b}}(\cdot)$. Thus
$m^{\prime}({\breve{G}}_{\lambda,\mu}(t_{2}))-m^{\prime}({\breve{G}}_{\lambda,\mu}(t_{1}))\geq 0\,$,
which implies that ${\breve{G}}_{\lambda,\mu}(t_{2})\geq{\breve{G}}_{\lambda,\mu}(t_{1})$, since $m^{\prime}(\cdot)$ is non-decreasing. Hence, ${\breve{G}}_{\lambda,\mu}$ is indeed non-decreasing.
Part 2. Left-continuity. We show that ${\breve{G}}_{\lambda,\mu}(\cdot)$ is left-continuous at the point $t\in\{s\in(0,1)\ |\ {\breve{G}}_{\lambda,\mu}(s)<+\infty\}$. Let $\{t_{n}\}_{n\in\mathbb{N}}$ be an non-decreasing sequence, $t_{n}\leq t_{n+1}$ for all $n\in\mathbb{N}$, that satisfies $\lim_{n\to\infty}t_{n}=t$. Next, we denote $g(n):={\breve{G}}_{\lambda,\mu}(t_{n})$, $n\in\mathbb{N}$, and $g(\infty):={\breve{G}}_{\lambda,\mu}(t)$. To obtain left-continuity we need to show that $g:=\lim_{n\to+\infty}g(n)=g(\infty)$. By non-decreasingness of the sequence $\{t_{n}\}_{n\in\mathbb{N}}$ and of ${\breve{G}}_{\lambda,\mu}(\cdot)$, it holds that $g(n)\leq g(\infty)$ for all $n\in\mathbb{N}$. Thus, by to the Monotone Convergence Theorem
$g=\lim_{n\to+\infty}g(n)\leq g(\infty)\,.$
As the sequence $g(\cdot)$ is non-decreasing, have that $g(n)\leq g$, for all $n\in\mathbb{N}$. By the definition of ${\breve{G}}_{\lambda,\mu}(t)$, see (12) and (10), we have
$g(n)=$ $\inf\left\{y\geq 0\ |\ \tfrac{\partial}{\partial y}\,h_{t_{n}}(y)\geq 0\right\}\,$
$=$ $\inf\left\{y\geq 0\ |\ m^{\prime}(y)+\lambda{\breve{F}}_{\varphi_{T}}(1-t_{n})-\mu\phi^{\prime}\big({\breve{F}_{b}}(t_{n})\big)\geq 0\right\}\,.$
As $\frac{\partial}{\partial y}h_{t_{n}}(y)$ is strictly increasing in $y$, we obtain that for all $\delta>0$ and $n\in\mathbb{N}$,
(16) $m^{\prime}(g(n)-\delta)<-\lambda{\breve{F}}_{\varphi_{T}}(1-t_{n})+\mu\phi^{\prime}\big({\breve{F}_{b}}(t_{n})\big)\leq m^{\prime}(g(n)+\delta).$
Consider the case when ${\breve{G}}_{\lambda,\mu}(t)$ is not left-continuous at $t$, that is when $g<g(\infty)$. Then take $\delta=\frac{g(\infty)-g}{2}>0$, which implies that
$g(n)+\delta=g(n)+\frac{g(\infty)-g}{2}\leq\frac{g(\infty)+g}{2}=g(\infty)-\frac{g(\infty)-g}{2}=g(\infty)-\delta\,.$
Thus, from (16) we derive that
(17a) $-\lambda{\breve{F}}_{\varphi_{T}}(1-t_{n})+\mu\phi^{\prime}\big({\breve{F}_{b}}(t_{n})\big)$ $\leq m^{\prime}(g(n)+\delta)$
(17b) $\leq m^{\prime}(g(\infty)-\delta)$
(17c) $<-\lambda{\breve{F}}_{\varphi_{T}}(1-t)+\mu\phi^{\prime}\big({\breve{F}_{b}}(t)\big)\,.$
Since ${\breve{F}}_{\varphi_{T}}(\cdot)$ is continuous and ${\breve{F}_{b}}(\cdot)$ is left-continuous, it holds that
$-\lambda{\breve{F}}_{\varphi_{T}}(1-t)+\mu\phi^{\prime}\big({\breve{F}_{b}}(t)\big)=$ $\lim_{n\to\infty}-\lambda{\breve{F}}_{\varphi_{T}}(1-t_{n})+\mu\phi^{\prime}\big({\breve{F}_{b}}(t_{n})\big)$
$<$ $-\lambda{\breve{F}}_{\varphi_{T}}(1-t)+\mu\phi^{\prime}\big({\breve{F}_{b}}(t)\big)\,$,
where the strict inequality follows from (17). Thus, we arrived at a contradiction to $g<g(\infty)$, which implies that ${\breve{G}}_{\lambda,\mu}(\cdot)$ is indeed left-continuous at $t$.
Finally, we establish the optimal quantile function to problem ($\breve{P}$). In particular, the optimal quantile function is given in (12) with a particular choice of Lagrange multipliers.
Theorem 3.11 (Optimal quantile function).
Let Assumptions 2.1 and 3.1 be satisfied. Then there exists a unique quantile function that solves ($\breve{P}$), which is given by
${\breve{F}}^{*}(t)={\breve{G}}_{\lambda^{*},\mu^{*}}(t),$
with ${\breve{G}}_{\lambda^{*},\mu^{*}}(t)$ defined in (12) and the Lagrange multipliers $\lambda^{*},\mu^{*}\geq 0$ satisfy the slackness conditions
$\lambda^{*}\left({\mathfrak{c}}\big({\breve{F}}^{*}\big)-x_{0}\right)$ $=0\,,\quad\text{and}$
$\mu^{*}\left({\mathscr{B}}_{\phi}\big({\breve{F}}^{*},{\breve{F}_{b}}\big)-{\varepsilon}\right)$ $=0\,.$
Proof 3.12.
By Theorems 3.5 and A.1, there exists a pair of multipliers $\lambda^{*}$ and $\mu^{*}$ such that the Karush-Kuhn-Tucker (KKT) conditions hold. Denote by $\breve{F}^{*}\in\mathcal{Q}_{\varepsilon}(x_{0})$ the solution satisfying the KKT conditions and recall that $\breve{F}^{*}$ is also the solution to (8) with Lagrange multipliers $(\lambda^{*},\mu^{*})$. We proceed to show that $\breve{F}^{*}(t)={\breve{G}}_{\lambda^{*},\mu^{*}}(t)$ for all $t\in(0,1)$, where ${\breve{G}}_{\lambda^{*},\mu^{*}}$ is the solution to (12) with Lagrange multipliers $(\lambda^{*},\mu^{*})$.
Denote by $\mathcal{I}_{1}:=\{t\in(0,1)\ |\ {\breve{G}}_{\lambda^{*},\mu^{*}}(t)<+\infty\}$ and by $\mathcal{I}_{2}:=\{t\in(0,1)\ |\ {\breve{G}}_{\lambda^{*},\mu^{*}}(t)=+\infty\}$. We show by contradiction that $\mathcal{I}_{2}=\emptyset$. For this assume that $\mathcal{I}_{2}\neq\emptyset$. Since $\breve{G}_{\lambda^{*},\mu^{*}}(t)$ is non-decreasing (see Proposition 3.9), it follows that $\mathcal{I}_{2}$ must be an interval of the form $(a,1)$ or $[a,1)$ for some $a\in[0,1)$. Consequently, we have
$\int_{0}^{1}{\breve{G}}_{\lambda^{*},\mu^{*}}(t){\mathrm{d}}t=\int_{\mathcal{I}_{1}}{\breve{G}}_{\lambda^{*},\mu^{*}}(t){\mathrm{d}}t+\int_{\mathcal{I}_{2}}{\breve{G}}_{\lambda^{*},\mu^{*}}(t){\mathrm{d}}t=+\infty\,.$
However, under 3.1, the feasible solution $\breve{G}_{\lambda^{*},\mu^{*}}(t)$ satisfies
$\int_{0}^{1}\breve{G}_{\lambda^{*},\mu^{*}}(t)\,\mathrm{d}t<+\infty\,$,
(see Proposition 3.2). This leads to a contradiction. Therefore, we have $\mathcal{I}_{2}=\emptyset$ and $\mathcal{I}_{1}=(0,1)$, which implies that $\breve{G}_{\lambda^{*},\mu^{*}}(t)$ is finitely-valued on $(0,1)$. By Proposition 3.7, ${\breve{G}}_{\lambda^{*},\mu^{*}}$ uniquely minimizes the Lagrangian over the set of functions mapping $(0,1)$ to $0,\infty)$. Since ${\breve{G}}_{\lambda^{*},\mu^{*}}$ is a quantile function itself, it also minimize the Lagrangian over the set of quantile functions and we conclude that $\breve{F}^{*}(t)={\breve{G}}_{\lambda^{*},\mu^{*}}(t)$ for all $t\in(0,1)$.
Next, we show that the optimal quantile function has a simple representation.
Corollary 3.13.
Let Assumptions 2.1 and 3.1 be satisfied. Then the unique quantile function that solves ($\breve{P}$) has representation
${\breve{F}}^{*}(t)=(m^{\prime})^{-1}\Big(-\lambda^{*}{\breve{F}}_{\varphi_{T}}(1-t)+\mu^{*}\phi^{\prime}\big({\breve{F}_{b}}(t)\big)\Big)\,$,
where $(m^{\prime})^{-1}$ is the inverse of $m^{\prime}(y)=-u^{\prime}(y)+\mu^{*}\phi^{\prime}(y)$, and $\lambda^{*}$ and $\mu^{*}$ are given in Theorem 3.11.
Proof 3.14.
From the proof of Theorem 3.11, it holds that the optimal quantile function is finite, i.e. ${\breve{F}}^{*}(t)={\breve{G}}_{\lambda^{*},\mu^{*}}(t)<+\infty$, for all $t\in(0,1)$. Thus, from Proposition 3.7 we have that for each $t\in(0,1)$, ${\breve{G}}_{\lambda^{*},\mu^{*}}(t)$ is the solution to
(18) $\frac{\partial}{\partial y}\;h^{*}_{t}(y)=m^{\prime}(y)+\lambda^{*}{\breve{F}}_{\varphi_{T}}(1-t)-\mu^{*}\phi^{\prime}\big({\breve{F}_{b}}(t)\big)=0\,$,
where $h^{*}_{t}(\cdot)$ is given in (10) with multipliers $\lambda^{*}$ and $\mu^{*}$. Solving (18) for $y$ concludes the proof.
From Theorems 3.5 and 3.11 we observe that there are three possible cases for the optimal quantile function:
  1. $i)$
    Both constraints are binding, i.e., $\lambda^{*},\mu^{*}>0$. Then, ${\breve{F}}^{*}(t)={\breve{G}}_{\lambda^{*},\mu^{*}}(t)$ and
    ${\mathfrak{c}}\big({\breve{F}}^{*}\big)=x_{0}\quad\text{and}\quad{\mathscr{B}}_{\phi}\big({\breve{F}}^{*},{\breve{F}_{b}}\big)={\varepsilon}\,.$
  2. ii)ii)
    Only the budget constraint is binding, i.e., $\lambda^{*}>0$ and $\mu^{*}=0$. Then, ${\breve{F}}^{*}(t)={\breve{G}}_{\lambda^{*},0}(t)$ and
    ${\mathfrak{c}}\big({\breve{F}}^{*}\big)=x_{0}\,.$
  3. $iii)$
    Only the BW constraint is binding, i.e., $\lambda^{*}=0$ and $\mu^{*}>0$. Then, ${\breve{F}}^{*}(t)={\breve{G}}_{0,\mu^{*}}(t)$ and
    ${\mathscr{B}}_{\phi}\big({\breve{F}}^{*},{\breve{F}_{b}}\big)={\varepsilon}\,.$
4 Examples
In this section we illustrate the theoretical results. We derive optimal payoffs and evaluate how these depend on the choice of the BW divergence, that is, on how investors reward positive deviations and penalize negative ones. We first layout the market model and then discuss how to choose the tolerance level $\varepsilon$} and the Bregman generating function $\phi$.
4.1 Market model
We provide examples within the Geometric Brownian Motion (GBM) market framework. In this market, there is a stock $(S_{t})_{t\in[0,T]}$ and a risk-free asset $(B_{t})_{t\in[0,T]}$ satisfying the stochastic differential equations
${\mathrm{d}}S_{t}=\mu_{s}S_{t}{\mathrm{d}}t+\sigma_{s}S_{t}{\mathrm{d}}W_{t},\quad\text{and}\quad{\mathrm{d}}B_{t}=rB_{t}{\mathrm{d}}t,$
respectively, where $(W_{t})_{t\in[0,T]}$ is a standard Brownian motion, and the stock has drift $\mu_{s}\in\mathbb{R}$ and instantaneous volatility $\sigma_{s}>0$. Under the no-arbitrage condition the state price density $\varphi_{T}$ is uniquely given by
$\varphi_{T}=e^{-\left(r+\frac{1}{2}\theta^{2}\right)T-\theta W_{T}},$
in which $\theta=\frac{\mu_{s}-r}{\sigma_{s}}$. Hence, $\varphi_{T}$ follows a log-normal distribution with mean $\mu_{\varphi}:=-(r+\frac{1}{2}\theta^{2})T$ and standard deviation $\sigma_{\varphi}:=\theta\sqrt{T}$, denoted as $\varphi_{T}\sim\text{LN}(\mu_{\varphi},\sigma_{\varphi}^{2})$. Its cdf and quantile function are
$F_{\varphi_{T}}(x)=\Phi\left(\frac{\ln x-\mu_{\varphi}}{\sigma_{\varphi}}\right)\quad\text{and}\quad\breve{F}_{\varphi_{T}}(t)=e^{\mu_{\varphi}+\sigma_{\varphi}\breve{\Phi}(t)},$
where $\Phi(\cdot)$ and $\breve{\Phi}(\cdot)$ are the cdf and quantile function of the standard normal distribution, respectively.
In all our numerical experiments, we assume the following parameters for the market model
$T=5\,,\quad r=0\,,\quad\mu_{s}=0.05\,,\quad\sigma_{s}=0.1\,,\quad\text{and}\quad S_{0}=1\,.$
Furthermore, we assume that the investor makes their choices using a Constant Relative Risk Aversion (CRRA) utility function
[…]
in which the parameter $\gamma$ reflects their degree of risk aversion. In the numerical experiments we choose the parameter values $\gamma=1$ and $\gamma=1.5$. Moreover, without any loss of generality, we assume that the investor has a budget $x_{0}=1$.
4.2 Choice of tolerance level $\varepsilon$}
When determining their optimal payoff as a solution to problem (P), the investor needs to choose the benchmark quantile function $\breve{F}_{b}$}, the tolerance level $\varepsilon$}, and the Bregman generating function $\phi$ that determines how deviations from the benchmark $\breve{F}_{b}$} are assessed.
To this regard, we consider the smallest BW distance ${\varepsilon}_{\min}$ for which problem ($\breve{P}$) is well-defined in that any $\varepsilon$} choice with ${\varepsilon}>{\varepsilon}_{\min}$ implies that 2.1 is satisfied, and together with 3.1 we obtain that a solution to ($\breve{P}$) exists. Specifically, we consider the problem
(19) $\underset{c(X_{T})\leq x_{0}}{\min}{\mathscr{B}}_{\phi}(\breve{F}_{X_{T}},{\breve{F}_{b}})\,$,
with the value ${\varepsilon}_{\min}$. Since ${\mathscr{B}}_{\phi}(\breve{F}_{X_{T}},{\breve{F}_{b}})$ only depends on the distribution of $X_{T}$, and cost-efficiency yields the least expensive way to achieve the distribution (c.f. Bernard et al., (2014), Corollary 2), it is sufficient to consider the case where $X_{T}$ is anti-monotonic with $\varphi_{T}$ under the budget $x_{0}$. That is
(20) $\underset{{\breve{G}}\in\mathcal{Q},{\mathfrak{c}}({\breve{G}})\leq x_{0}}{\min}{\mathscr{B}}_{\phi}({\breve{G}},{\breve{F}_{b}})\,.$
In applications, it is often reasonable that the initial budget $x_{0}$ satisfies $x_{0}={\mathfrak{c}}\big({\breve{F}}_{b}\big)$, which implies that ${\varepsilon}_{\min}=0$, i.e., the investor can choose any ${\varepsilon}>0$ as tolerance level for deviations from the benchmark and ($\breve{P}$) is still well-posed. The following proposition provides for any given budget $x_{0}$ and benchmark $\breve{F}_{b}$} the expression for ${\varepsilon}_{\min}$.
Proposition 4.1.
For a budget $x_{0}$ and a benchmark $\breve{F}_{b}$}, the minimal BW distance ${\varepsilon}_{\min}$ is given by
[…]
where
${\breve{G}}^{*}(t):=\left(\phi^{\prime}\right)^{-1}\left(\phi^{\prime}\left({\breve{F}}_{0}(t)-\eta^{*}{\breve{F}}_{\varphi_{T}}(1-t)\right)\right),$
and $\eta^{*}>0$ is the unique solution to ${\mathfrak{c}}\big({\breve{G}}^{*}\big)=x_{0}$.
Proof 4.2.
If ${\mathfrak{c}}({\breve{F}_{b}})\leq x_{0}$, then ${\varepsilon}_{\min}=0$. To see this note that the Bregman divergence $B_{\phi}(z_{1},z_{2})$ is zero, if and only if $z_{1}=z_{2}$, thus ${\mathscr{B}}_{\phi}({\breve{G}},{\breve{F}})=0$ if and only if ${\breve{G}}(u)={\breve{F}}(u)$ for all $u\in(0,1)$.
For the second case, assume that ${\mathfrak{c}}({\breve{F}_{b}})>x_{0}$. Then the associated Lagrangian to (20) with Lagrange multiplier $\eta\geq 0$ is
$\int_{0}^{1}\phi\big({\breve{G}}(t)\big)-\phi\big({\breve{F}_{b}}(t)\big)-\phi^{\prime}\big({\breve{F}_{b}}(t)\big)\Big({\breve{G}}(t)-{\breve{F}_{b}}(t)\Big){\mathrm{d}}t+\eta\Big(\int_{0}^{1}{\breve{G}}(t){\breve{F}}_{\varphi_{T}}(1-t){\mathrm{d}}t-x_{0}\Big)\,.$
The remainder of the proof follows using similar arguments as in Propositions 3.7, 3.9, and Theorem 3.11.
For a given tolerance level $\varepsilon$}, we observe from Figure 1 that the Bregman divergence between two points vary significantly for different $\phi$. Therefore, we propose to make $\varepsilon$} dependent on the choice of $\phi$. To this end, in a similar spirit to Bernard et al., 2024b , the investor considers a series of strategies that they deem acceptable, and then select the tolerance level $\varepsilon$} as the maximum of the divergences of the acceptable strategies with the benchmark.
4.3 Choice of Bregman generator $\phi$
With the BW divergence, an investor can penalize underperformance and outperformance relative to the benchmark asymmetrically. Moreover, as investors may not want to penalize gains when measured relative to the terminal wealth of the benchmark, see e.g., Harlow, (1991) and Klebaner et al., (2017), we introduce a family of Bregman generators who assess deviations in relation to a wealth threshold $\alpha\in[0,\infty)$.
Definition 4.3 (Bregman divergence with threshold $\alpha$).
Let $\phi\colon\mathbb{R}\to\mathbb{R}$ be a convex and continuously differentiable Bregman generator. For a given threshold $\alpha$, we define the Bregman generator $\tilde{\phi}(\cdot;\alpha)$ associated with $\phi$ as
(21) $\tilde{\phi}(x;\alpha)=\phi(x)\,\mathbb{I}_{\{x\leq\alpha\}}+\left(\phi^{\prime}(\alpha)(x-\alpha)+\phi(\alpha)\right)\,\mathbb{I}_{\{x>\alpha\}},$
where $\mathbb{I}_{\{\cdot\}}$ denotes the indicator function. Then the Bregman divergence with generator $\tilde{\phi}(\cdot;\alpha)$ becomes
$B_{\tilde{\phi}}(z_{1},z_{2};\alpha)$ $=B_{\phi}(z_{1},z_{2})\,\mathbb{I}_{\{z_{1},z_{2}\leq\alpha\}}+B_{\phi}(z_{1},\alpha)\,\mathbb{I}_{\{z_{1}\leq\alpha<z_{2}\}}$
$\quad+\left\{B_{\phi}(\alpha,z_{2})+\big(\phi^{\prime}(\alpha)-\phi^{\prime}(z_{2})\big)(z_{1}-\alpha)\right\}\,\mathbb{I}_{\{z_{2}\leq\alpha<z_{1}\}}\,.$
Note that $\tilde{\phi}$ is continuously differentiable and convex, thus all relevant results of problem (P) apply to the Bregman divergence with threshold $\alpha$.
We observe that if $z_{1},z_{2}\leq\alpha$, then $B_{\tilde{\phi}}$ reduces to the Bregman divergence with $\phi$ and if $z_{1},z_{2}>\alpha$, then $B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha\big)=0$, meaning that the Bregman divergence no longer accounts for the distance between points that both exceed the threshold $\alpha$. If $z_{1}\leq\alpha<z_{2}$, the divergence equals $B_{\phi}(z_{1},\alpha)$, that is only deviation of $z_{1}$ to the reference point $\alpha$ are taken into account, independent of the value of $z_{2}$. Moreover, due to the asymmetry of the Bregman divergence, if $z_{2}\leq\alpha<z_{1}$, the penalization is not only $B_{\phi}(\alpha,z_{2})$ but also via the derivative of $\phi$. The Bregman divergence with threshold $\alpha$ thus combines the tractability of Bregman divergence with the ability to measure downside risk relative to the wealth threshold $\alpha$ and to moreover penalize underperformance relative to the benchmark. To gain a deeper understanding of the role of the threshold $\alpha$, Proposition 4.4 establishes the monotonicity and limit properties of ${\mathscr{B}}_{\tilde{\phi}}({\breve{F}}_{1},{\breve{F}}_{2};\alpha)$ with respect to $\alpha$. Its proof is relegated to Appendix B.
Proposition 4.4.
For ${\breve{F}}_{1},{\breve{F}}_{2}\in{\mathcal{Q}}$, the function $\alpha\mapsto{\mathscr{B}}_{\tilde{\phi}}({\breve{F}}_{1},{\breve{F}}_{2};\alpha)$ is non-decreasing. Moreover, it holds that
(22) $\lim_{\alpha\to 0^{+}}{\mathscr{B}}_{\tilde{\phi}}({\breve{F}}_{1},{\breve{F}}_{2};\alpha)=0,\quad\text{and}\quad\lim_{\alpha\to\infty}{\mathscr{B}}_{\tilde{\phi}}({\breve{F}}_{1},{\breve{F}}_{2};\alpha)={\mathscr{B}}_{\phi}({\breve{F}}_{1},{\breve{F}}_{2}).$
The key emphasis of the divergence ${\mathscr{B}}_{\tilde{\phi}}({\breve{F}}_{1},{\breve{F}}_{2};\alpha)$ is when ${\breve{F}}_{1}$ and ${\breve{F}}_{2}$ do not simultaneously exceed the threshold $\alpha$. Specifically, for $t\in(0,1)$, if both ${\breve{F}}_{1}(t)$ and ${\breve{F}}_{2}(t)$ are less than $\alpha$, then $B_{\tilde{\phi}}\big({\breve{F}}_{1}(t),{\breve{F}}_{2}(t);\alpha\big)=B_{\phi}\big({\breve{F}}_{1}(t),{\breve{F}}_{2}(t)\big)$, indicating that the divergence is measured using the Bregman divergence with $\phi$. However, if ${\breve{F}}_{1}(t)$ and ${\breve{F}}_{2}(t)$ lie on opposite sides of $\alpha$, then $B_{\tilde{\phi}}\big({\breve{F}}_{1}(t),{\breve{F}}_{2}(t);\alpha\big)\leq B_{\phi}\big({\breve{F}}_{1}(t),{\breve{F}}_{2}(t)\big)$, thereby penalizing less than with $B_{\phi}$, and if both ${\breve{F}}_{1}(t)$ and ${\breve{F}}_{2}(t)$ are larger than the wealth threshold $\alpha$ the divergence is zero.
Alternatively, we can view ${\mathscr{B}}_{\tilde{\phi}}$ as a function of $\alpha$. If $\alpha=0$, the investor’s problem (P) reduces to the case without a divergence constraint, that is the classical Merton problem. As $\alpha$ increases to infinity, ${\mathscr{B}}_{\tilde{\phi}}$ converges to the BW divergence with $\phi$. Therefore, $\alpha$ can be regarded as a weighting parameter that interpolates between these two extreme cases.
In the numerical examples below, we use the following two convex functions as Bregman generators
$\phi_{1}(x)=x^{2}\quad\text{and}\quad\phi_{2}(x)=x\ln x,\quad\text{for}\quad x>0\,.$
The choice $\phi_{1}$ corresponds to the 2-Wasserstein distance where positive and negative deviations from the benchmark quantiles are penalized equally. By contrast, under $\phi_{2}$ deviations are penalized in an asymmetric manner in that negative deviations (underperforming the benchmark) receive more weight than positive ones (outperforming the benchmark). Further, note that the convexity of $\phi_{2}$ becomes less strong when $x$ increases, thus underperforming the benchmark is penalised significantly more. We further use these two convex functions as generators for the Bregman divergence with given threshold $\alpha$, which then become
$\tilde{\phi}_{1}(x;\alpha)$ $=x^{2}\mathbb{I}_{\{x\leq\alpha\}}+2\alpha(x-\frac{\alpha}{2})\mathbb{I}_{\{x>\alpha\}}\,$,
$\tilde{\phi}_{2}(x;\alpha)$ $=x\ln x\mathbb{I}_{\{x\leq\alpha\}}+\big((\ln\alpha+1)x-\alpha\big)\mathbb{I}_{\{x>\alpha\}}\,.$
Note that both generators are linear in $x$ on ($\alpha$$\infty$).
Following Proposition 3.2 and Theorem 3.5, to ensure the existence of an optimal solution, we adopt the approach described in Remark 3.4. Specifically, in the numerical examples below, we add a regularization term of the form $\frac{\vartheta}{2}x^{2}$ to $\phi_{2}(x)$, $\tilde{\phi}_{1}(x;\alpha)$, and $\tilde{\phi}_{2}(x;\alpha)$. Unless otherwise stated, we set $\vartheta=10^{-8}$ by default. For notational simplicity, the original symbols are used to denote their regularized versions.
4.4 Example 1: Optimal payoff when the benchmark is a constant
We study optimal payoff choice for an investor who aims to maximize his expected utility under the constraint that their terminal wealth does not diverge too much from the terminal wealth when pursuing an investment bearing a fixed rate of return $\kappa>0$. That is, we assume in problem (P) that the benchmark’s quantile function is ${\breve{F}_{b}}\equiv e^{\kappa T}$. This situation corresponds to the case of pension fund managers whose performance and solvency position might be assessed with reference to such a benchmark. For instance, under a fair value approach, the supervisory authorities may require that ${\breve{F}_{b}}\equiv e^{\kappa T}$, where $\kappa$ is equal to the risk free rate $r$.
In order to determine the appropriate tolerance level $\varepsilon$}, the investor considers three investment strategies that he deems acceptable (although not necessarily optimal). For each of these strategies, the investor calculates their BW divergence to the benchmark’s and defines $\varepsilon$} as the maximum among them (one for each of the Bregman generator $\phi$ we study). The three acceptable strategies are as follows:
Strategy 1: The constant mix strategy is a strategy where one maintains, through continuous rebalancing, a portfolio which during the entire investment period $[0,T]$ keeps the proportions invested in the riskless asset respectively in the stock constant. We consider $82.5\%$ invested in the risk-free asset and the remaining proportion of $17.5\%$ is invested in stock, in which case, the quantile function of the terminal wealth is
${\breve{F}}_{T}(u)=e^{\mu_{1}+\sigma_{1}\breve{\Phi}(u)},\quad u\in(0,1)\,$,
where $\mu_{1}:=\left(r+\left(\mu_{s}-r\right)0.175-\frac{1}{2}0.175^{2}{\sigma^{2}_{s}}\right)T$ and $\sigma_{1}:=0.175\,\sigma_{s}\,\sqrt{T}$.
Strategy 2: For this strategy one invests at $t=0$ a proportion of the initial wealth in the risk-free asset and the remaining proportion in the stock. This is a buy and hold strategy, meaning that during $(0,T]$ there are no trades, thus the relative exposures to the risk-free asset and to the stock evolves over time. We consider $85\%$ of the initial wealth invested in the risk-free asset and the remaining proportion of $15\%$ is invested in stock. The terminal wealth’s quantile function is
${\breve{F}}_{T}(u)=\left(0.15e^{\mu_{2}+\sigma_{2}\breve{\Phi}(u)}+0.85e^{rT}\right),\quad u\in(0,1)\,$,
where $\mu_{2}:=\left(\mu_{s}-\frac{1}{2}\sigma^{2}_{s}\right)T$ and $\sigma_{2}:=\sigma_{s}\sqrt{T}$.
Strategy 3: This strategy yields a so-called digital payoff $Y_{T}$ on the stock $S_{T}$, given by
$Y_{T}=0.9\,\mathbb{I}_{\{S_{T}\leq c\}}+\tfrac{1}{0.95}\,(e^{rT}-0.045)\,\mathbb{I}_{\{S_{T}>c\}}\,$,
where $c>0$ is such that $\mathbb{Q}(S_{T}<c)=0.05$. Note that the cost of $Y_{T}$ is indeed equal to $x_{0}=1$. Its quantile function is given as
${\breve{F}}_{T}(u)=0.9\,\mathbb{I}_{\{u\leq\alpha_{c}\}}+\tfrac{1}{0.95}(e^{rT}-0.045)\,\mathbb{I}_{\{u>\alpha_{c}\}},\quad u\in(0,1)\,$,
where $\alpha_{c}:=\mathbb{P}(S_{T}<c)$.
Figure 2: The quantile functions of the three acceptable strategies (left panel) and the relationship between their payoffs and the stock price (right panel).
Figure 2 illustrates the quantile functions of the three acceptable strategies (left panel) and the relationship between their payoffs and the stock price (right panel). Notably, the quantile functions of strategies 1 and 2 have similar shapes, however their lower bounds ${\breve{F}}_{T}(0^{+})$ differ significantly. For strategy 1 the lower bound is 0, whereas for strategy 2 the lower bound is $0.85$. Thus, strategy 1 carries more risk in the event of extreme stock price losses, see also the right panel. As compared with the first two strategies, strategy 3 provides the best protection against downside risk, which however comes at the cost of limited upside potential.
Table 1: BW divergences between each acceptable strategy’s terminal wealth and that of the benchmark’s. The top row corresponds to the BW divergence with generator $\phi_{1}$, while the second row to the generator $\phi_{2}$. The last column provides the chosen tolerance level $\varepsilon$}, that is the maximal BW divergences of each row.
BW generator BW divergence  chosen $\varepsilon$}
Strategy 1 Strategy 2 Strategy 3
$\phi_{1}$ $0.003673$ $0.003717$ $0.000526$ $0.003717$
$\phi_{2}$ $0.001785$ $0.001799$ $0.000272$ $0.001799$
Table 1 reports for each admissible strategy the BW divergence of its terminal wealth to that of the benchmark’s. The last column displays the maximal $\varepsilon$} for each row, which is the tolerance level the investor chooses. The first row correponds to the Bregman generator $\phi_{1}$ and the second to the generator $\phi_{2}$.
We first study the optimal payoff when the investor does not care about the benchmark, that is when the BW constraint is absent and $\varepsilon$$\infty$ in problem (P). Figure 3 displays the optimal quantile functions for risk aversion parameters $\gamma=1$ and $\gamma=1.5$. Note that when $\gamma=1.5$ the quantile function is, as expected, flatter. We observe that these quantile functions are far away from the benchmark.
Figure 3: The optimal quantile function ${\breve{F}}^{*}(\cdot)$ in the absence of a BW constraint, i.e., $\varepsilon$$\infty$ in problem ($\breve{P}$), for $\gamma=1$ (blue line) and $\gamma=1.5$ (red line), respectively. The green curve depicts the quantiles of the constant benchmark.
Next, Figure 4 displays the optimal quantile function for problem ($\breve{P}$) when the BW constraint is applied. We observe that the shape of the optimal quantile functions changes in a significant manner and the variation in $\gamma$ has a smaller impact on $\breve{F}$$^{*}$, indicating that the BW constraint becomes the dominant factor. We make the following additional observations. First, the similarity of $\breve{F}$$^{*}$ under the different $\phi_{i}$ suggests that our method for choosing $\varepsilon$} is not unreasonable. Second, in both the left and right panels we see that - to satisfy the BW constraint - the optimal payoffs are less likely to fall below the benchmark payoff and tend to stay near it. Specifically, the optimal payoff has about 55% probability of falling below that of the benchmark’s, while the upper bound of the payoff, ${\breve{F}}^{*}(1^{-})$, is around $1.073$.
Figure 4: The optimal quantile functions ${\breve{F}}^{*}(u)$ for the generators $\phi_{1}(x)=x^{2}$ (blue lines), $\phi_{2}(x)=x\ln x$ (red lines), and the constant benchmark ${\breve{F}_{b}}(u)$ (green line) for $\gamma=1$ and $\gamma=1.5$, respectively.
To further explore the behaviour of the optimal payoffs, Figure 5 displays how they depend on the stock $S_{T}$. In this regard, recall that $\hat{X}_{T}={{{\breve{F}}^{*}}}(1-F_{\varphi_{T}}(\varphi_{T}))$. Since $\mu_{s}\ >r$, it holds that $\varphi_{T}$ is monotonically decreasing in $S_{T}$, which results that the optimal payoffs are non-decreasing in $S_{T}$. For $\gamma=1$, we observe that when the stock price declines by more than 60 percent, i.e., when $S_{T}<0.4$, then all optimal payoffs are close to zero. For $0.4\leq S_{T}<1$, the optimal payoffs under BW constraints increase very rapidly with $S_{T}$ and clearly surpass the optimal payoff without BW constraints. Moreover, we find that under BW constraints, optimal payoffs will outperform the benchmark as soon as $S_{T}>0.85$, also reflecting a capacity for risk mitigation. However, their outperformance will always remain fairly moderate. In contrast, the optimal payoff without the BW constraint only begins to outperform the benchmark when $S_{T}>1.05$, at which point the outperformance increases rapidly. Similar findings hold for the case in which $\gamma=1.5$ (right panel of Figure 5).
Figure 5: The optimal payoff $\hat{X}_{T}$ as a function of $S_{T}$ under the BW constraint induced by the generators $\phi_{1}(x)=x^{2}$ (blue lines), $\phi_{2}(x)=x\ln x$ (red lines) and when no such constraint is considered (purple lines). The left panel depicts the optimal payoffs for $\gamma=1$ and the right panel for $\gamma=1.5$.
Next, we study the optimal payoff under the BW divergence that are driven by generators $\tilde{\phi}_{1}(\cdot;\alpha)$ and $\tilde{\phi}_{2}(\cdot;\alpha)$ with threshold values $\alpha=0.95$ and $\alpha=1$. Table 2 provides, similarly to Table 1, the BW divergence values for the three acceptable strategies and the different BW divergences. Again, the last column depicts the tolerance levels $\varepsilon$} that the investor chooses.
Table 2: BW divergence between each acceptable strategy’s terminal wealth and that of the benchmark’s. The different rows corresponds to different BW divergences with generators indicated in the first column. The last column provides the chosen tolerance level $\varepsilon$}, that is the maximal BW divergences of each row.
BW generator BW divergence   chosen $\varepsilon$}
Strategy 1 Strategy 2 Strategy 3
$\tilde{\phi}_{1}(\cdot;1)$ $0.000088$ $0.000065$ $0.000500$ $0.000500$
$\tilde{\phi}_{2}(\cdot;1)$ $0.000045$ $0.000033$ $0.000259$ $0.000259$
$\tilde{\phi}_{1}(\cdot;0.95)$ $0.000002$ 0.0 $0.000125$ $0.000125$
$\tilde{\phi}_{2}(\cdot;0.95)$ $0.000001$ 0.0 $0.000067$ $0.000067$
Figure 6: Optimal quantile functions ${\breve{F}}^{*}(u)$ for the generators $\tilde{\phi}_{1}(\cdot;\alpha)$ (blue lines), $\tilde{\phi}_{2}(\cdot;\alpha)$ (red lines) and the constant benchmark ${\breve{F}_{b}}(u)$ (green line). The BW generator threshold is $\alpha=0.95$ (bottom panels) and $\alpha=1$ (top panels) and the investor’s risk aversion is $\gamma=1$ (left panels) and $\gamma=1.5$ (right panels), respectively.
Figure 7: Optimal quantile functions ${\breve{F}}^{*}(u)$ for the generators $\tilde{\phi}_{1}(\cdot;1)$ (blue lines), $\tilde{\phi}_{2}(\cdot;1)$ (red lines) and the constant benchmark ${\breve{F}_{b}}(u)$ (green line) with different choice of $\vartheta$. The investor’s risk aversion is $\gamma=1$.
Figure 8: Optimal payoffs $\hat{X}_{T}$ as a function of $S_{T}$ for the generators $\tilde{\phi}_{1}(\cdot;\alpha)$ (blue lines) and $\tilde{\phi}_{2}(\cdot;\alpha)$ (red lines). The purple line shows the payoff when the BW divergence constraint is absent. The BW generator threshold is $\alpha=0.95$ (bottom panels) and $\alpha=1$ (top panels) and the investor’s risk aversion is $\gamma=1$ (left panels) and $\gamma=1.5$ (right panels), respectively.
Figure 6 presents the optimal quantile functions for the BW divergence with threshold $\alpha$. We observe that when ${\breve{F}}^{*}(u)$ is below the threshold value $\alpha=1$ in the top panels ($\alpha=0.95$ in the bottom panels), the optimal quantile function and the benchmark quantile function are close. This is in contrast to when ${\breve{F}}^{*}(u)$ exceeds $\alpha$, in which case the optimal quantile function increases rapidly. This indicates that the divergence constraint primarily targets the downside risk and penalizes outcomes that fall below the threshold $\alpha$. The magnified insets in each subplot illustrate the details of the quantile function ${\breve{F}}^{*}(u)$ within the interval $[0,0.005]$ (bottom left) and when it crosses the threshold $\alpha$ (top right). Notably, as $\gamma$ increases (from left to right panels), ${\breve{F}}^{*}(u)$ reaches $\alpha$ earlier, but the subsequent rate of increase becomes slower. This indicates that risk averse investors (larger $\gamma$) are more inclined to secure the return level $\alpha$ rather than to pursue higher returns. In summary, since the budget and the $\varepsilon$} are fixed, if investors wish to reduce downside risk, they must correspondingly forgo high returns during favourable market conditions.
Figure 7 showcases the effect of different regularization parameters $\vartheta$ on the numerical computation of the optimal quantile functions. It can be observed that the overall shapes of the optimal quantile functions remain similar across different values of $\vartheta$. As $\vartheta$ decreases, the influence of the regularization term becomes weaker. In fact, when $\vartheta=10^{-6}$, the resulting optimal quantile functions are already very close to those shown in the upper-left panel of Figure 6 ($\vartheta=10^{-8}$). In practical applications, a smaller value of $\vartheta$ can be chosen to achieve higher numerical accuracy.
Figure 8 illustrates the relationship between the various payoffs $\hat{X}_{T}$ and the stock price $S_{T}$. As compared to Figure 5, we observe that the payoffs are no longer subject to an upper bound, allowing investors to achieve higher returns when the stock price is elevated. Furthermore, the optimal payoffs present strong risk resilience during periods of declining stock prices. The different figures clearly show that using the BW constraint with a wealth threshold makes it possible to effectively balance downside risk (by staying close to the constant benchmark) against the potential to take advantage of more favourable market returns.
4.5 Example 2: Optimal payoff under a non-constant benchmark
In this section, we consider the case of an investor with a more aggressive benchmark strategy. Specifically, the benchmark strategy has a constant exposure of $80\%$ in the stock and $20\%$ in the risk-free asset, that is maintained during the investment horizon. The quantile function of terminal wealth the benchmark strategy generates is
${\breve{F}_{b}}(u)=e^{(r+(\mu_{s}-r)0.8-\frac{1}{2}0.8^{2}\,\sigma_{s}^{2})T+0.8\,\sigma_{s}\,\sqrt{T}\,\breve{\Phi}(u)},\ u\in(0,1).$
To determine the tolerance level $\varepsilon$}, the investor pursues similarly to Section 4.4 by considering the BW divergence to acceptable strategies. Since the benchmark strategy is more aggressive to the constant benchmark in Section 4.4, we modify the three acceptable strategies to reflect the investor’s choice of benchmark.
Strategy 1: The constant mix strategy with 75% invested in the stock and the remaining 25% in the risk-free asset.
Strategy 2: The buy-and-hold strategy with 85% invested in the stock and the remaining 15% in the risk-free asset.
Strategy 3: A digital payoff $Y_{T}$ on the stock $S_{T}$ given by
$Y_{T}=0.8\,\mathbb{I}_{\{S_{T}\leq c\}}+\tfrac{1}{0.9}(e^{rT}-0.08)\,\mathbb{I}_{\{S_{T}>c\}}.$
where $c>0$ is such that $\mathbb{Q}(S_{T}<c)=0.1$.
As before, we consider Bregman generators $\phi_{i}(\cdot)$, $\tilde{\phi}_{i}(\cdot;1)$ and $\tilde{\phi}_{i}(\cdot;0.95)$ for $i=1,2$. Table 3 presents the BW divergences between each acceptable strategy’s terminal wealth and that of the benchmark as well as the investor’s chosen tolerance level $\varepsilon$} for each of the Bregman generators.
Table 3: Choice of the tolerance levels $\varepsilon$}. The first three columns display the divergences for the three strategies for all different generators that we consider. The last column provides the resulting tolerance level $\varepsilon$}.
BW generator BW divergence  chosen $\varepsilon$}
Strategy 1 Strategy 2 Strategy 3
$\phi_{1}(\cdot)$ 0.000506 0.001179 0.086821 0.086821
$\phi_{2}(\cdot)$ 0.001785 0.000367 0.032795 0.032795
$\tilde{\phi}_{1}(\cdot;1)$ 0.000007 0.000009 0.001108 0.001108
$\tilde{\phi}_{2}(\cdot;1)$ 0.000004 0.000005 0.000630 0.000630
$\tilde{\phi}_{1}(\cdot;0.95)$ 0.000007 0.000007 0.001023 0.001023
$\tilde{\phi}_{2}(\cdot;0.95)$ 0.000004 0.000004 0.000586 0.000586
Figure 9: The optimal payoff as a function of $S_{T}$ when there are no BW constraints (purple curve) and when there is a BW constraint induced by the generator $\phi_{1}(x)=x^{2}$ (blue curve) or by the generator $\phi_{2}(x)=x\ln x$ (red curve). The non constant benchmark is depicted by the green curve. The cases $\gamma=1$ (left panel) and $\gamma=1.5$ (right panel) are studied.
Figure 10: The optimal payoff as a function of $S_{T}$, when there are no BW constraints (purple curve) and when there is a BW constraint induced by the generator $\tilde{\phi}_{1}(x;\alpha)$ (blue curve) or by the generator $\tilde{\phi}_{2}(x;\alpha)$ (red curve). The non constant benchmark is depicted by the green curve. The cases $\gamma=1$ (left panel) and $\gamma=1.5$ (right panel) are studied.
Figure 9 and Figure 10 illustrate the relationship between the optimal payoffs $\hat{X}_{T}$ and the stock price $S_{T}$. We observe that, overall, the payoff of the benchmark increases in a linear fashion with $S_{T}$. This is due to the large allocation of stock within the benchmark. Furthermore, when the BW constraint is absent, $\varepsilon$$\infty$, (purple curve) the optimal payoff is very different from the Benchmark. Adding the BW constraints brings, as expected, the optimal payoffs (blue and red curves) closer to the benchmark. When comparing the impact of the BW generators $\phi_{i}$, $i=1,2$, (Figure 9) with the BW generators with threshold $\tilde{\phi_{i}}$, $i=1,2$, (Figure 10), in the latter case the optimal payoffs stay closer to the benchmark when the underlying stock is declining and significantly outperform the benchmark when the stock is appreciating.
In summary, the BW constraints allows the investor to create payoffs that yield large expected utility while staying close to the benchmark and in particular to obtain payoffs that show resilience when markets are declining.
5 Concluding remarks
In this paper, we deal with optimal payoff choice for an expected utility maximizer who aims to stay close to a benchmark, but considers asymmetry in that positive deviations are measured differently than negative ones. Doing so has significant impact on the payoff choice. By acknowledging and accounting for the asymmetric preferences of investors, financial professionals can design products and strategies that better align with their clients’ preferences and risk tolerances.
Acknowledgement
SP acknowledges support from the Natural Sciences and Engineering Research Council of Canada (grants DGECR-2020-00333 and RGPIN-2020-04289). SV is grateful to FWO for financial support (grant numbers FWO SBO S006721N and FWO WOG W001021N). JY acknowledges the support from the National Natural Science Foundation of China (12371474).
Appendix A Convex Optimization in Banach Spaces
In this appendix, we collect auxiliary results on convex optimisation problems.
Theorem A.1.
Consider $L$ as a real linear space, and let $f:L\to\bar{\mathbb{R}}$ be a convex function. Define the feasible set $\mathcal{F}_{L}$ as follows:
$\mathcal{F}_{L}:=\left\{x\in L\ |\ g_{i}(x)\leq 0,\ i=1,\ldots,n,\ r_{j}(x)=0,\ j=1,\ldots,m\right\},$
where $g_{i}$ and $r_{i}$ are $\bar{\mathbb{R}}$-valued functions on $L$ which are convex and affine, respectively. Assume that $\mathcal{F}_{L}$ is nonempty and consider the optimization problem
$\min\{f(x)\ |\ x\in\mathcal{F}_{L}\}.$
If $x^{*}\in\mathcal{F}_{L}$ is an solution to this problem, i.e.,
$f(x^{*})=\min\{f(x)\ |\ x\in\mathcal{F}_{L}\},$
then there exist $n+m+1$ real numbers $\nu^{*},\lambda_{1}^{*},\ldots,\lambda_{n}^{*},\mu_{1}^{*},\ldots,\mu_{m}^{*}$ satisfying:
  • 1.
    $\nu^{*}\geq 0$;
  • 2.
    $\lambda_{i}^{*},\mu_{j}^{*}\geq 0$ for all $i=1,\ldots,n,j=1,\ldots m$;
  • 3.
    $\nu^{*}f(x^{*})\leq\nu^{*}f(x)+\sum_{i=1}^{n}\lambda_{i}^{*}g_{i}(x)+\sum_{j}^{m}\mu_{j}^{*}r_{j}(x)$ for all $x\in L$;
  • 4.
    $\lambda_{i}^{*}g_{i}(x^{*})=0$ for all $i=1,\ldots,n$.
Conditions 1 – 4 are commonly referred to as the Karush–Kuhn–Tucker (KKT) conditions, with condition 4 specifically known as the complementary slackness condition.
Proof A.2.
See Theorem 3.1 in Barbu and Precupanu, (2012).
The numbers $\lambda_{i}^{*}$ and $\mu_{j}^{*}$, possessing the properties outlined in the theorem, are referred to as the Lagrange multipliers. We observe that it suffices to consider $\nu^{*}=0$ or $\nu^{*}=1$. To see this, note that if $\nu^{*}>0$ and $\nu\neq 1$, we can redefine $\lambda_{i}^{*}:=\lambda_{i}^{*}/\nu^{*}$, $i=1,\ldots,n$, and $\mu_{j}^{*}:=\mu_{j}^{*}/\nu^{*}$, $j=1,\ldots,m$, which still satisfy the conditions in Theorem 2.3.
Definition A.3 (Lagrangian function).
For the above convex optimization problem, define its Lagrangian function as
$\mathcal{L}(x,\lambda,\mu)=f(x)+\sum_{i=1}^{n}\lambda_{i}g_{i}(x)+\sum_{j=1}^{m}\mu_{j}r_{j}(x),$
for all $(x,\lambda,\mu)\in L\times[0,\infty)^{n}\times\mathbb{R}^{m}$.
Theorem A.1 provides the necessary conditions for a solution to exist. With certain additional assumptions, we can demonstrate that these conditions are also sufficient.
Definition A.4 (Slater condition).
An optimization problem satisfies the Slater condition if there exists a point $x_{0}\in\mathcal{F}_{L}$ such that
$g_{i}(x_{0})<0,\ \ \forall i=1,\ldots,n.$
Theorem A.5.
Under the assumptions of Theorem A.1, if we additionally assume that the Slater condition holds, then the necessary conditions of Theorem A.1 become sufficient for optimality, which implies $\nu^{*}>0$. Furthermore, it is sufficient to optimize the Lagrangian function $\mathcal{L}(x,\lambda^{*},\mu^{*})$ with respect to $x$ over $L$.
Proof A.6.
See Theorems 3.7 - 3.9 in Barbu and Precupanu, (2012).
Appendix B Proof of Proposition 4.4
This appendix is devoted to additional proofs not provided in the main body of the paper.
Proof B.1.
We first prove that $\alpha\mapsto B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha\big)$ is non-decreasing for fixed non-negative $z_{1}$ and $z_{2}$. First, let $\alpha_{1}\leq z_{1}\leq\alpha_{2}\leq z_{2}\leq\alpha_{3}$, then we prove that
$B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{1}\big)\leq B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{2}\big)\leq B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{3}\big).$
From (21), we have
$B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{1}\big)=$ $\ 0,$
$B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{2}\big)=$ $\ \phi(z_{1})-\phi(\alpha_{2})-\phi^{\prime}(\alpha_{2})(z_{1}-\alpha_{2}),$
$B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{3}\big)=$ $\ \phi(z_{1})-\phi(z_{2})-\phi^{\prime}(z_{2})(z_{1}-z_{2})=B_{\phi}\big(z_{1},z_{2}\big).$
Note that only $B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{2}\big)$ depends on the threshold $\alpha$. Let $\alpha_{2,1},\alpha_{2,2}\in[z_{1},z_{2}]$, with $\alpha_{2,1}\leq\alpha_{2,2}$. Then using in the second equation the Mean Value Theorem with $\xi\in[\alpha_{2,1},\alpha_{2,2}]$, we obtain
$B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{2,2}\big)-B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{2,1}\big)$
$=$ $-\phi(\alpha_{2,2})+\phi(\alpha_{2,1})-\phi^{\prime}(\alpha_{2,2})(z_{1}-\alpha_{2,2})+\phi^{\prime}(\alpha_{2,1})(z_{1}-\alpha_{2,1})$
$=$ $-\phi^{\prime}(\xi)(\alpha_{2,2}-\alpha_{2,1})-\left(\phi^{\prime}(\alpha_{2,2})-\phi^{\prime}(\alpha_{2,1})\right)z_{1}+\phi^{\prime}(\alpha_{2,2})\alpha_{2,2}-\phi^{\prime}(\alpha_{2,1})\alpha_{2,1}$
$=$ $\left(\phi^{\prime}(\alpha_{2,2})-\phi^{\prime}(\xi)\right)\alpha_{2,2}-\left(\phi^{\prime}(\alpha_{2,2})-\phi^{\prime}(\alpha_{2,1})\right)z_{1}+\left(\phi^{\prime}(\xi)-\phi^{\prime}(\alpha_{2,1})\right)\alpha_{2,1}\,.$
Since $\phi^{\prime}(\cdot)$ is non-decreasing, we have $\phi^{\prime}(\alpha_{2,2})-\phi^{\prime}(\xi)\geq\phi^{\prime}(\alpha_{2,2})-\phi^{\prime}(\alpha_{2,1})$. Combing with $\alpha_{2,2}\geq z_{1}$ and that $\alpha_{2,1}\geq 0$, we get
(23) $B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{2,2}\big)-B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{2,1}\big)\geq\left(\phi^{\prime}(\xi)-\phi^{\prime}(\alpha_{2,1})\right)\alpha_{2,1}\geq 0,$
which implies that $\alpha_{2}\mapsto B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{2}\big)$ is non-decreasing on $z_{1},z_{2}]$. Thus, we obtain
(24) $B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{1}\big)=B_{\tilde{\phi}}\big(z_{1},z_{2};z_{1}\big)\leq B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{2}\big)\leq B_{\tilde{\phi}}\big(z_{1},z_{2};z_{2}\big)=B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha_{3}\big).$
Similarly, when $z_{1}>z_{2}$, we can prove that the inequalities (23) and (24) hold, i.e., $\alpha\mapsto B_{\tilde{\phi}}\big(z_{1},z_{2};\alpha\big)$ is non-decreasing. Thus, we conclude that $\alpha\mapsto{\mathscr{B}}_{\tilde{\phi}}({\breve{F}}_{1},{\breve{F}}_{2};\alpha)$ is non-decreasing. Finally, taking the limit in $\alpha$, we obtain (22).